### v2

# Finite Population Identification and

Abstract
We develop a new approach for quantifying uncertainty in finite populations, by using design distributions to calibrate sensitivity parameters in finite population identified sets. This yields uncertainty intervals that can be interpreted as identified sets, Bayesian credible sets, or frequentist design-based confidence sets. We focus on quantifying uncertainty about the average treatment effect (ATE) due to missing potential outcomes in a randomized experiment, where our approach (1) yields design-based confidence intervals for ATE which allow for heterogeneous treatment effects but do not rely on asymptotics, (2) provides a new motivation for examining covariate balance, and (3) gives a new formal analysis of the role of randomized treatment assignment. We illustrate our approach in three empirical applications.
JEL classification: C18; C21; C25; C51
Keywords: Treatment Effects, Partial Identification, Randomization, Uncertainty Quantification
1 Introduction
Since the beginning of statistics, there have been two ways to think about populations:
  1. 1.
    They are finite, as in Neyman (1923), or Bowley (1926) who considered “a sample containing $n$ persons or things [that] is selected from a universe containing $N$….” (page 3).
  2. 2.
    Or they are infinite, as in Fisher (1930), or Student (1908) who assumed “a random sample has been obtained from an indefinitely large* population…. *Note that the indefinitely large population need not actually exist” (page 302).
Inference in finite populations came to be called “design-based inference” or “randomization inference” while inference in infinite populations came to be called “model-based inference” or “super-population inference” (e.g., Wu and Thompson 2020). Econometric theory has traditionally focused on super-population inference, going back to Haavelmo (1944) who explained that “when we describe $s$ as a random variable with a certain probability distribution for each fixed set of values of the variables $x$, we are thinking of a class of hypothetical, infinite populations…” (page 51). Likewise, much of statistics has also focused on the super-population approach, with survey sampling theory as the main exception (see Little 2004). [^3] These distinctions and jargon come from the survey sampling literature (e.g., Särndal et al. 1978), which is how we use them here. In economics, some researchers now use the term “design-based” in a related but not equivalent sense, to refer to specific identification strategies (e.g. Card 2022). We focus on randomized experiments in this paper and hence both meanings of “design-based” apply.
However, in the past decade there has been a resurgence of interest in design-based inference in explicitly finite populations (see section 1.2). Much of this work follows a template largely laid out by Neyman (1923). This template has three steps:
  1. 1.
    Describe the finite population.
  2. 2.
    Describe the design, which explains how the data is obtained.
  3. 3.
    Derive properties of the induced design distribution of statistics that depend on the population and the data.
For example, in the context of survey sampling, the population is a set of numbers $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$ (step 1). A classical design is simple random sampling: Sample exactly $n$ units from $\{1,\ldots,N\}$ uniformly at random (step 2). Let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be binary random variables indicating which units were sampled. A classical theorem is that the sample mean (a function of $\mathbf{S}$ and $\mathbf{Y}$) is design-unbiased for the finite population mean when $\mathbf{S}$ is chosen uniformly at random: $\mathbb{E}_{\text{design}}\left[\frac{1}{n}\sum_{i=1}^{N}S_{i}Y_{i}\right]=\frac{1}{N}\sum_{i=1}^{N}Y_{i}$ (step 3). Viewed from the perspective of econometrics, there is a noticeable omission in this template: What about identification? Identification analysis is a key step throughout much of super-population based econometrics (e.g., Lewbel 2019 or Molinari 2020); is any of that analysis still useful in finite populations? Yes, as we argue in this paper. We use ideas from identification theory to construct a new approach for quantifying uncertainty in finite populations, which we call design-based sensitivity analysis. This approach is non-asymptotic, which allows it to be reliably applied in small populations. It yields uncertainty intervals that can be interpreted as (1) identified sets, (2) Bayesian credible sets, or (3) frequentist design-based confidence sets. And it allows researchers to flexibly impose additional assumptions on unobserved values, transparently quantifying the impact of these assumptions on conclusions.
We develop this approach in the context of randomized experiments with a binary treatment, where the goal is to quantify the uncertainty in the average treatment effect (ATE) due to missing potential outcomes (we also briefly discuss sampling uncertainty). In this context, our approach yields three further contributions: (1) Our uncertainty intervals allow for heterogeneous treatment effects, implying that our approach yields design-based confidence intervals for ATE without relying on asymptotics or imposing homogeneous treatment effects. These intervals are not constructed via hypothesis test inversion, but rather are based on ideas from partial identification. (2) We give a new motivation for examining covariate balance. Specifically, by making an assumption that explicitly links covariates to potential outcomes, we show that observed imbalances in covariates have identifying power for the unobserved realized imbalance in potential outcomes, which allows researchers to obtain tighter bounds on ATE. And (3) we give a new formal analysis of the role of randomized treatment assignment.
1.1 Our approach to quantifying uncertainty in experiments: A practitioner’s guide
Here we briefly sketch our approach, with an eye toward practical take-aways. Consider the population of six units in table 1 (left). Here $Y_{i}(1)$ and $Y_{i}(0)$ denote potential outcomes and $X_{i}\in\{0,1\}$ denotes realized treatment. The true ATE is 0.55, the difference between the average of $Y_{i}(1)$ over $i=1,\ldots,6$ and the average of $Y_{i}(0)$ over $i=1,\ldots,6$. The data, shown in the right table, only reveals half of the potential outcomes, however. Suppose we know that all outcomes must be between 0 and 1. Then without further assumptions, all we can say for sure about ATE is that it is in the set
$\left[\frac{(0.6+0.9+0.8)+(0+0+0)}{6}-\frac{(0.1+0+0.2)+(1+1+1)}{6},\right.$ (1)
$\quad\left.\frac{(0.6+0.9+0.8)+(1+1+1)}{6}-\frac{(0.1+0+0.2)+(0+0+0)}{6}\right]=\left[-\frac{1}{6},\frac{5}{6}\right].$
This is the finite population “worst case” identified set for ATE (following Manski’s 1990 derivations), obtained by filling in the unobserved values in table 1 (right) with values between $[0,1]$ to either maximize or minimize the corresponding ATE value. However, suppose we also knew that treatment was randomly assigned. Unfortunately, as is well known (e.g., Altman 1985, Greenland and Robins 1986, and Greenland 1990), randomization does not guarantee any kind of ex post balance in potential outcomes. Hence knowledge that treatment was randomized does not shrink the finite population identified set (we formalize this in theorem 2). But what if we were willing to assume that the average of the unobserved values in the dashed box are not farther than $K$ from the observed mean $(0.1+0+0.2)/3=0.1$ in the solid box? And likewise suppose we are willing to assume that the average of the unobserved values in the dashed oval are not farther than $K$ from the observed mean $(0.6+0.9+0.8)/3=0.77$ in the solid oval (a finite population version of an assumption proposed in Manski 2003). Then for sufficiently small $K$, we can conclude that ATE is in a strictly smaller set than equation (1), where the size of this set depends on just how small $K$ is. That is, its length depends on just how balanced we think the average potential outcomes are between the top 3 and bottom 3 rows of table 1 (right). Step 1 of our approach is to plot these identified sets as a function of $K$, the maximal magnitude of allowed imbalance in average potential outcomes. The top plot in figure 1(a) shows an example (using real data; see section 8.1). By examining how these sets change with $K$, we can examine the sensitivity of conclusions about ATE to assumptions about the magnitude of the realized differences in average potential outcomes across the treatment and control group.
$i$ $Y_{i}(0)$ $Y_{i}(1)$ $X_{i}$
1 0.1 0.8 0
2 0 0.5 0
3 0.2 0.7 0
4 0.4 0.6 1
5 0.1 0.9 1
6 0.2 0.8 1
$i$ $Y_{i}(0)$ $Y_{i}(1)$ $X_{i}$
1 0.1 ? 0
2 0 ? 0
3 0.2 ? 0
4 ? 0.6 1
5 ? 0.9 1
6 ? 0.8 1
Table 1: Left: Example population, $N=6$. Right: The observed data from that population.
The challenge for any sensitivity analysis, including this one, is to provide a meaningful interpretation of the magnitude of the sensitivity parameter. In this case, which values of $K$ should be considered “large” and which should be considered “small”? To answer this, in step 2 we convert the values of $K$ into probabilities. Specifically, suppose the true values of all potential outcomes were known. Then, since we know the treatment assignment distribution, we can compute the ex ante probability that the difference in potential outcome means across the treatment and control groups is at most $K$. This can be done for any $K$. Since we do not actually know all potential outcomes, we can find the smallest possible value of this probability, across all logically possible completions of table 1 (right). Denote this smallest probability by $\underline{p}(K)$. Step 2 of our procedure is to plot this function $\underline{p}(\cdot)$, which converts $K$-values into ex ante design-probabilities. The bottom plot in figure 1(a) shows an example.
(a) Top: $\Theta_{I}(K)$, bounds on ATE as a function of the maximum magnitude of ex post imbalance, $K$. Bottom: $\underline{p}(K)$.
(b) $\Theta_{I}(K(\alpha))$, bounds on the ATE as a function of $1-\alpha$.
Figure 1: Example output for a design-based sensitivity analysis, $N=10$. See section 8.1 for data details.
Finally, in step 3, we recommend combining these two plots as follows: For any desired ex ante probability $1-\alpha\in(0,1)$, use the bottom plot to find the smallest magnitude of imbalance that occurs with at least $100(1-\alpha)$ex ante probability; denote this value by $K(\alpha)$. For example, if we set $0.91-\alpha=$ 0.9 then the figure shows us that $K=18.9$ satisfies $\underline{p}(K)=1-\alpha$ (see the dotted lines). We can then move to the top plot to read off bounds on ATE for this value of the sensitivity parameter, as shown in the dotted lines. In this case the bounds are about $[-6,22]$. Figure 1(b) shows these bounds for all values of $1-\alpha$. We show that these uncertainty intervals have three interpretations:
  1. 1.
    First, they can be interpreted as finite population identified sets for ATE, under various deterministic assumptions on the magnitude of ex post imbalance.
  2. 2.
    Second, they can be interpreted as robust Bayesian credible sets, which contain the true ATE with probability at least $1-\alpha$, according to the “robust empirical objective” prior $\underline{p}$. There is a true, but unknown, magnitude of imbalance in potential outcomes (e.g., it is 0.13 and 0.1 in table 1). Formally, we show that $\underline{p}$ is a valid cumulative distribution function, and thus can be used to model one’s beliefs about the largest value of the magnitude of this ex post imbalance. Furthermore, any prior on potential outcomes implies a prior on this true magnitude of ex post imbalance. We show that $\underline{p}$ stochastically dominates all such implied priors, implying that is a “worst case” or “robust” prior.
  3. 3.
    Third, for any fixed $1-\alpha$, we show that these bounds are design-based confidence intervals. That is, across repeated re-assignments of treatment, these bounds will contain the true ATE at least $100(1-\alpha)$of the time.
Overall, we recommend that researchers report plots like figure 1(b), which can then be interpreted any of these ways, depending on personal preference.
Researchers also often report summary statistics that are meant to measure the “strength of evidence” in favor of a hypothesis. For example, we may ask: How much evidence is there in the data for the conclusion that ATE is nonnegative? The horizontal intercept in figure 1(b), called a breakdown point, provides a simple answer. In figure 1(a), we first compute the horizontal intercept in the top plot, which is the largest magnitude of imbalance between mean potential outcomes that can be allowed while still allowing us to conclude that ATE is nonnegative. We then plug it into $\underline{p}$ to convert it to design probabilities (see the dashed lines). Here we obtain $\underline{p}(12.871)=0.65$. This tells us that, given the data, and regardless of what the true unknown potential outcomes actually are, there was an ex ante probability of at least 65% that the treatment and control groups would be sufficiently balanced to ensure that the identified set for ATE only contains non-negative numbers. Using the Bayesian interpretation of our analysis, this number says there is at least a 65% chance that ATE is non-negative, according to our randomization based prior distribution.
In some applications, the bounds we just described, as in figure 1(b), may be deemed large for many values of $\alpha$. In this case, we can obtain tighter bounds by making further assumptions about unobserved potential outcomes. In many experiments researchers observe covariates for each unit. It is often reasonable to expect that these covariates are predictive of potential outcomes. If researchers are willing to place a lower bound on this predictive power, as measured by the R-squared from the regression of potential outcomes onto covariates, then they can obtain tighter uncertainty intervals for ATE. We illustrate this analysis in figure 4(b) in section 8.2.
Beyond this basic setup, we discuss a variety of extensions and complementary results throughout the paper, including (1) instrumental variables and noncompliance, (2) random sampling of units, in addition to randomization, (3) other restrictions on the unobserved potential outcomes, like having an a priori bounded variance, (4) identification of parameters besides the average treatment effect, and (5) using other measures of balance beyond means.
1.2 Related literature
Our paper relates to a variety of literatures, but for brevity we only discuss the most closely related work, rather than attempt a comprehensive survey. Manski and Pepper (2018) derive identified sets in an explicitly finite population setting. Like them, we also consider the identifying power of bounded variation type assumptions, which restrict just how far potential outcomes can be from each other. The main difference is that they focus on a setting with observational data whereas we consider randomized experiments. Our focus on experiments allows us to use randomization itself to calibrate the magnitude of the bounded-variation sensitivity parameter, which is the design-based sensitivity analysis we develop in this paper. Like us, Greenland and Robins (1986) also describe the problem of drawing conclusions in finite populations as an identification problem which must be solved by making assumptions about unobserved values of variables. They formally show how exact balance assumptions yield point identification. They then informally explain that “if we randomize…when both samples are large, random differences will in probability be small” (page 415). Many other papers make similar informal remarks, like Lindley (1980, page 590) and Royall and Pfeffermann (1982, page 480). Relatedly, the distinction between ex ante and ex post balance is well known; Greenland (1990) gives a particularly clear discussion, where he considers a randomized experiment with $N=2$. Greenland and Robins (2009) more recently discuss this distinction, and conclude that “randomization (or more generally, ignorability) does not impose “no [ex post] confounding”…rather, it provides…a randomization-based (“objective”) derivation of a prior…that applies after allocation as well as before, and becomes more narrowly centered around zero as the sample size increases. This is a key post-allocation benefit of randomization” (page 5, emphasis in the original). That paper as well as the rest of this prior literature, however, is largely verbal discussion. One of our main contributions is to use those informal ideas to provide a new, detailed, and formal approach to quantifying uncertainty in small finite populations.
The standard approach to uncertainty quantification in finite populations is to construct frequentist design-based confidences sets. For example, chapter 5 of Imbens and Rubin (2015) discusses confidence sets which are valid for finite $N$ but are based on homogeneous treatment effect assumptions, while chapter 6 discusses confidence sets which allow for heterogeneous treatment effects but which are based on asymptotics. Our uncertainty intervals also have a confidence set interpretation, and so fit within this standard approach. Importantly, however, our intervals also have two additional non-frequentist interpretations, as either identified sets or Bayesian credible sets. We are not aware of any other intervals in the literature which have these three distinct interpretations. Moreover, viewed purely as confidence sets for ATE, our intervals have valid coverage for any fixed $N$, allow for heterogeneous treatment effects, and do not require outcomes to be binary. We are only aware of one feasible alternative design-based confidence interval with these properties, which we call the Hoeffding CI, as discussed in Aronow et al. (2025) and Ding (2025) (who build on the binary outcome analysis of Robins and Ritov 1997). In our empirical applications this interval is always at least 3.5 times wider than ours, and can be up to 20 times wider, depending on the nominal coverage probability (see appendix LABEL:sec:AlternativeEmpiricalCIs). Since both intervals have valid coverage, this suggests that the Hoeffding CI is substantially less powerful than our interval. We discuss these issues as well as further related literature in appendix LABEL:sec:simulations.
Finally, many recent papers on causal inference study explicitly finite populations, including Li and Ding (2017), Ding et al. (2017), Aronow and Samii (2017), Ding (2017), Athey et al. (2018), Kang et al. (2018), Abadie et al. (2020, 2023), Hong et al. (2020), Rambachan and Roth (2020), Wu and Ding (2021), Eckles et al. (2020), Imbens and Menzel (2021), Zhao and Ding (2021), Sävje (2021), Bojinov et al. (2021), Xu (2021), Xu and Wooldridge (2022), Athey and Imbens (2022), Roth and Sant’Anna (2023), Pollmann (2023), Wooldridge (2023), Startz and Steigerwald (2023), Borusyak et al. (2024), and Sancibrián (2024), among many others. Also see Imbens and Rubin (2015) and Ding (2024) for book level surveys. This literature largely follows the template of Neyman (1923) that was discussed earlier, and does not do explicit identification analysis.
2 Finite Population Identification
We first define the finite population identified set. We then use this definition to derive identified sets for finite population average treatment effects under various assumptions.
2.1 The finite population identified set
Consider the standard potential outcomes model with a binary treatment. There are four components required to define an identified set:
1. The population: Suppose there are $N$ units in our population. Let =$\mathcal{I}=\{1,\ldots,N\}$ be the set of indices for these units. Each unit $i\in\mathcal{I}$ is associated with the vector of numbers $(Y_{i}(1),Y_{i}(0),X_{i},W_{i})$, where $Y_{i}(1)$ and $Y_{i}(0)$ are potential outcomes, $X_{i}\in\{0,1\}$ is a realized binary treatment, and $W_{i}$ is a $d_{W}$-vector of covariates. The population is the $N\times(3+d_{W})$ matrix of all of these numbers. Let $\mathbf{P}^{\text{true}}\coloneqq(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{X},\mathbf{W})$ denote this matrix. We use $\mathbf{P}$ to denote alternative possible values of this population matrix. Table 1 shows an example population.
2. Assumptions: We do not observe $\mathbf{P}^{\text{true}}$. We will make assumptions about it, however. Formalize these assumptions as the restriction that $\mathbf{P}^{\text{true}}\in\mathcal{P}$ where $\mathcal{P}$ is a known set of $N\times(3+d_{W})$ matrices. We give examples of $\mathcal{P}$ in section 2.2.
3. Parameters: Parameters of interest are functionals of the population matrix. For example, the average treatment effect is defined as $\text{ATE}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\big{(}Y_{i}(1)-Y_{i}(0)\big{)}$. Similarly, the average effect of treatment on the treated is $\text{ATT}\coloneqq\Big{(}\sum_{i=1}^{N}\big{(}Y_{i}(1)-Y_{i}(0)\big{)}\allowbreak\mathbbm{1}(X_{i}=1)\Big{)}/\sum_{i=1}^{N}\mathbbm{1}(X_{i}=1)$, assuming the denominator is nonzero. In general, let $\theta(\mathbf{P})$ be a functional defined on the set of logically possible values of the population matrix. Let $\Theta$ denote the set of logically possible values of this parameter. Let $\theta^{\text{true}}\coloneqq\theta(\mathbf{P}^{\text{true}})$ denote its true value.
4. The data: Finally, we must describe the data that is observed to the econometrician. To allow for sampling, let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be a vector of indicators where $S_{i}=1$ if unit $i$ appears in the dataset, and zero otherwise, as in section 1. Let $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$ where $Y_{i}\coloneqq Y_{i}(1)X_{i}+Y_{i}(0)(1-X_{i})$ for all $i\in\mathcal{I}$. Then $\mathbf{P}^{\text{data}}$ is the matrix of all rows of $(\mathbf{Y},\mathbf{X},\mathbf{W})$ with $S_{i}=1$. Formalize this construction via the function $\text{MakeData}(\mathbf{P},\mathbf{S})$, so that $\mathbf{P}^{\text{data}}=\text{MakeData}(\mathbf{P}^{\text{true}},\mathbf{S})$. [^4] We define the population at the time period after which treatment has been assigned but before units are sampled. This is standard in the identification literature (e.g., Manski 2003, ch. 7). One could modify the notation to further distinguish population values of potential outcomes and covariates from realized treatment assignment, as none of our concepts or procedures depend on this particular notational choice, but at the cost of additional notation throughout, especially since we mainly focus on the case where all units are sampled.
We are now ready to define the identified set.
Definition 1.
The identified set for $\theta$ is
$\Theta_{I}\coloneqq\{\theta\in\Theta:\theta=\theta(\mathbf{P})\text{ for some }\mathbf{P}\in\mathcal{P}\text{ such that }\text{MakeData}(\mathbf{P},\mathbf{S})=\mathbf{P}^{\text{data}}\}.$
This definition follows the usual description of the identified set as “the set of parameters…that are consistent with the model and the data” (Tamer 2010, page 184), but with two variations to clarify what is meant by “data” when the population is finite: First, even when all units are sampled, the data is defined by a finite dimensional matrix rather than a joint distribution of random variables (as in def. 3.1 on page 5324 of Matzkin 2007, for example; also see our appendix LABEL:sec:popDistributions for a related discussion). Second, our definition allows for data on some units to be missing (i.e., not sampled), by setting $S_{i}=0$. This is perhaps unusual, since uncertainty due to sampling is traditionally separated from other types of uncertainty, as in Koopmans’ (1949, page 132) original definition of identification. Nonetheless, as we show in this paper, definition 1 provides a useful foundation to quantify uncertainty from both missing units and missing potential outcomes. In particular, $\Theta_{I}$ has two key features: (1) It can always be computed with the data the analyst actually has at hand, and (2) like traditional definitions of identification, it is guaranteed to contain the true parameter of interest, so long as the model is not false. We will revisit these properties below.
Finally, note that some authors define the “data” as the design distribution of the observed data, rather than the dataset that is actually observed. This leads to a different definition of identification that can be applied to finite populations; for example, see the definition of “sampling identification” in Florens and Simoni (2021, def 2.1 on page 5). This design distribution typically depends on the entire population matrix and hence is never actually knowable, even if all units are sampled (because not all potential outcomes are observed). Moreover, under this alternative definition of identification, the existence of a design-unbiased estimator for a parameter implies that the parameter is point identified. In contrast, under definition 1, any parameter that depends on at least one unknown value, like a missing potential outcomes, will typically be partially identified, even if there is a design-unbiased estimator for it. In this paper we argue that this feature of our definition leads to a useful approach for quantifying uncertainty in finite populations.
2.2 Identified sets for ATE
To simplify the exposition, from here on we assume that all units in the population are observed, and focus solely on the uncertainty due to missing potential outcomes. In section 7.1 we show that the extension to incorporate sampling is straightforward. We also focus on the identification of the average treatment effect for brevity. We briefly discuss other parameters in section 7.3.
It is well known that, because averages are influenced by the values of outlier observations, bounds on ATE are usually infinite without some kind of restriction on the magnitude of outliers. Hence we maintain the following assumption throughout the paper.
Assumption A1 (Bounded outcomes).
There are known values $-\infty<y_{\text{min}}<y_{\text{max}}<\infty$ such that $Y_{i}(x)\in[y_{\text{min}},y_{\text{max}}]$ for all $i\in\mathcal{I}$, for each $x\in\{0,1\}$.
In some applications the values of $y_{\text{min}}$ and $y_{\text{max}}$ can be set to their logical values, such as 0 to 100 for test scores. In other settings, like when outcomes are wages, these values are sensitivity parameters that reflect our beliefs about the smallest and largest possible values of potential outcomes in the population under consideration. We discuss several variations on this assumption in section 7.2, including an assumption that does not require selecting a priori bounds on outcomes.
A1 alone has identifying power for ATE, via a finite population version of the Manski (1990) bounds. This result is a special case of a more general result that uses the following exogeneity assumption.
Assumption A2 ($K$-approximate mean balance).
There is a known $K\geq 0$ such that $K^{\text{true}}(x)\leq K$ for $x\in\{0,1\}$, where
$K^{\text{true}}(x)\coloneqq\left|\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)\right|$ (2)
and $N_{x}\coloneqq\sum_{i=1}^{N}\mathbbm{1}(X_{i}=x)$ is the number of units who receive treatment $x$.
This assumption was proposed by Manski (2003, page 149), who called it approximate mean independence. He noted that if $K=0$ then it is equivalent to mean independence of potential outcomes from realized treatment. Manski and Pepper (2018) call this a bounded variation type assumption. A2 could be generalized to allow a different $K$ for each potential outcome; we use a common $K$ for simplicity. Let $\overline{\mathbf{Y}(x)}\coloneqq\frac{1}{N}\sum_{i=1}^{N}Y_{i}(x)$ and $\overline{Y}_{x}\coloneqq\frac{1}{N_{x}}\sum_{i=1}^{N}Y_{i}\mathbbm{1}(X_{i}=x)$. With a minor variation on Manski’s (1990) derivations, we obtain the following characterization of the finite population identified set for ATE under the $K$-approximate mean balance assumption.
Theorem 1.
Suppose A1 and A2 hold, and $\mathbf{P}^{\text{data}}$ is known. Then, for each $x\in\{0,1\}$, the identified set for $\overline{\mathbf{Y}(x)}$ is $[\text{LB}_{K}(x),\text{UB}_{K}(x)]$ where
$\text{LB}_{K}(x)$ $\coloneqq\overline{Y}_{x}\frac{N_{x}}{N}+\max\{y_{\text{min}},\overline{Y}_{x}-K\}\frac{N_{1-x}}{N},\text{ and}$
$\text{UB}_{K}(x)$ $\coloneqq\overline{Y}_{x}\frac{N_{x}}{N}+\min\{y_{\text{max}},\overline{Y}_{x}+K\}\frac{N_{1-x}}{N}.$
Moreover, the identified set for ATE is $\Theta_{I}(K)\coloneqq[\text{LB}_{K}(1)-\text{UB}_{K}(0),\;\text{UB}_{K}(1)-\text{LB}_{K}(0)]$.
The magnitude of uncertainty about ATE depends on the choice of $K$. At one extreme, for sufficiently large values ($K\geq\overline{K}\coloneqq\max\{\overline{K}_{1},\overline{K}_{0}\}$ where $\overline{K}_{x}\coloneqq\max\{\overline{Y}_{x}+y_{\text{max}},\;\overline{Y}_{x}-y_{\text{min}}\}$), $\Theta_{I}(K)$ gives what Manski (2003) calls a “domain of consensus”, a set of ATE values that are consistent with the observed data and A1, but otherwise no restriction at all on the relationship between realized treatment and potential outcomes. Denote this set by $\Theta_{I}(\infty)$. At the other extreme, $K=0$, mean independence holds. In this case, ATE is point identified and equals the observed difference in means, $\overline{Y}_{1}-\overline{Y}_{0}$. In general, smaller $K$’s lead to smaller identified sets. So how should researchers assess the credibility of a choice of $K$?
2.3 Randomization and identification in finite populations
[…]
$\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)=\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)$
for $x\in\{0,1\}$. In fact, for many values of the vectors $\mathbf{Y}(x)=(Y_{1}(x),\ldots,Y_{N}(x))$, it is impossible for exact balance to hold regardless of the values of realized treatment $\mathbf{X}=(X_{1},\ldots,X_{N})$, a point noted by Greenland and Robins (1986, page 415). So statistical independence is not an appropriate formalization of random assignment in finite populations. Instead, we make the following assumption.
Assumption A3 (Random assignment).
The size of the treatment and control groups, $N_{1}$ and $N_{0}$, are fixed a priori. $\mathbf{X}=(X_{1},\ldots,X_{N})$ is a single realization from the known probability distribution $\mathbb{P}_{\text{design}}$ on $\{0,1\}^{N}$ defined by
$\mathbb{P}_{\text{design}}(X_{1}^{\text{new}}=x_{1},\ldots,X_{N}^{\text{new}}=x_{N})=1\Big{/}{N\choose N_{1}}$
for all $(x_{1},\ldots,x_{N})\in\{0,1\}^{N}$ with $\sum_{i=1}^{N}x_{i}=N_{1}$, and equal to 0 otherwise.
In this assumption we use the notation $\mathbf{X}^{\text{new}}=(X_{1}^{\text{new}},\ldots,X_{N}^{\text{new}})$ to denote the random vector with distribution $\mathbb{P}_{\text{design}}$. $\mathbf{X}=(X_{1},\ldots,X_{N})$ is a single, non-random realization of this random variable. Call $\mathbb{P}_{\text{design}}$ the design distribution of treatment. This particular choice of design distribution is often called uniform randomization. It is a standard formalization of randomization in the design-based inference literature; for example, see Imbens and Rubin (2015, section 4.4). We conjecture that most of our results extend to other design distributions commonly used in randomized experiments (as surveyed in Rosenberger and Lachin 2015, for example), but we focus on uniform randomization for brevity.
As has long been recognized, randomization in finite populations does not guarantee any level of ex post balance (e.g., Greenland 1990). The identified set, by definition, must contain the true parameter whenever the model is true. Consequently, for any finite population, randomization has no identifying power. The following theorem formalizes this result.
Theorem 2.
Suppose A1 and A3 hold and $\mathbf{P}^{\text{data}}$ is known. Then the identified set for ATE is $\Theta_{I}(\infty)$.
Put differently, after treatment has been assigned, we cannot rule out the possibility that potential outcomes are substantially imbalanced across the treatment and control groups, even if it was unlikely a priori. Thus, from an identification perspective, randomization does not shrink the domain-of-consensus bounds in finite populations. Nonetheless, in the next section we’ll show that theorem 2 provides an overly pessimistic view of the value of randomization. There we reinterpret randomization as a procedure that affects our beliefs about ex post balance. Specifically, we will use randomization to assess the plausibility of a specific choice of the sensitivity parameter $K$. This will let us use the identification result in theorem 1 to perform a sensitivity analysis motivated by random assignment of treatment.
3 Design-Based Sensitivity Analysis
3.1 A design-based approach to calibrating $K$
A traditional sensitivity analysis would plot the identified set $\Theta_{I}(K)$ as a function of the sensitivity parameter $K$, as in the top plot of figure 1(a). This plot shows how our conclusions can vary from point identification of ATE (under exact balance, $K=0$) to partial identification under approximate balance ($K>0$). Like any sensitivity analysis, however, there is an important question: Which values of $K$ are most plausible? In this section, we provide an objective approach to calibrating this sensitivity parameter, based on randomization. Specifically, define
$p(K,\mathbf{Y}(1),\mathbf{Y}(0))\coloneqq\\ \mathbb{P}_{\text{design}}\left(\left|\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}^{\text{new}}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}^{\text{new}}=0)\right|\leq K\text{ for $x\in\{0,1\}$}\right).$
This is the design-probability that the $K$-approximate mean balance assumption holds, when the true potential outcomes are $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$. Although $\mathbb{P}_{\text{design}}$ is known, $p(K,\mathbf{Y}(1),\mathbf{Y}(0))$ is unknown since it depends on the specific values of the potential outcomes. However, we can obtain bounds on this probability by using two constraints: (1) Half of all potential outcomes are known (namely, those associated with the observed treatments) and (2) Outcomes are bounded (A1). These restrictions combined with the known design distribution will let us obtain bounds on the probability $p(K,\mathbf{Y}(1),\mathbf{Y}(0))$ for all $K$. We specifically focus on lower bounds since these tell us the smallest design probability of $K$-approximate balance, which will lead to the most conservative inference. So let
$\underline{p}(K)\coloneqq\inf_{\begin{subarray}{c}\mathbf{Y}(1),\mathbf{Y}(0)\in[y_{\text{min}},y_{\text{max}}]^{N}\\ \mathbf{Y}=\mathbf{Y}(1)\times\mathbf{X}+\mathbf{Y}(0)\times(\mathbb{1}-\mathbf{X})\end{subarray}}\;p\big{(}K,\mathbf{Y}(1),\mathbf{Y}(0)\big{)}$ (3)
where $\times$ denotes component-wise multiplication and $\mathbb{1}$ is an $N$-vector of ones. $\underline{p}$ is a known function. It depends on the realized data $(\mathbf{Y},\mathbf{X})$ and the design distribution $\mathbb{P}_{\text{design}}$. We discuss how to feasibly compute this function in section 3.5. For now we focus on its interpretation and use. Note that, by definition, $\underline{p}(K)\leq p(K,\mathbf{Y}(1),\mathbf{Y}(0))$ for the true population values of $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$.
The key idea is to use $\underline{p}(K)$ to interpret the magnitude of the sensitivity parameter $K$ in the identification results of section 2. Specifically, we suggest performing what we call a design-based sensitivity analysis:
  1. 1.
    First, plot $\Theta_{I}(K)$, the identified set for ATE, as a function of $K$. Below this, plot the function $\underline{p}(K)$. Figure 1(a) gives an example of this paired plot. The top graph shows the sequence of finite population identified sets alone. The second plot then converts its horizontal axis values of $K$ into design-probabilities.
  2. 2.
    Focus on particular values of $K$. There are several reasonable ways to do this:
    1. (a)
      Define the breakdown point $K^{\text{bp}}\coloneqq\sup\{K\geq 0:0\notin\Theta_{I}(K)\}$, the largest relaxation of exact balance such that zero is not in the identified set. Researchers can compute $\underline{p}(K^{\text{bp}})$ to interpret this value. In figure 1(a) it is 0.65. Thus there was at least a 65% ex ante probability that potential outcomes would be sufficiently balanced that we can conclude that ATE is positive.
    2. (b)
      Next, notice that $\underline{p}(K)$ decreases as $K$ decreases, because smaller $K$ implies a stronger balance assumption that is therefore less likely to hold. Let $\alpha\in(0,1)$. Define $K(\alpha)\coloneqq\inf\{K\geq 0:\underline{p}(K)\geq 1-\alpha\}$, the closest we can get to exact balance while still ensuring that approximate balance holds with design-probability at least $1-\alpha$. Researchers can then present the set $\Theta_{I}\big{(}K(\alpha)\big{)}$ as a function of $1-\alpha$. Figure 1(b) gives an example.
3.2 Interpreting $\Theta_{I}(K(\alpha))$
Our main recommendation is that researchers report the set $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$, as in figure 1(b). This set can then be interpreted using any of three different interpretations. First, it can be interpreted as the finite population identified set for ATE under a sequence of deterministic assumptions on the magnitude of ex post imbalance.
Second, it has a robust Bayesian interpretation. Specifically, given randomization (A3), and for given prior beliefs about the potential outcomes, there is an implied prior distribution over the ex post difference in average potential outcomes between the treatment and control groups, $K^{\text{true}}(x)$ (defined in eq. (2)). However, it is unclear which prior on the potential outcomes to use. It turns out that $\underline{p}$ is a “worst case” or “robust” prior on $K^{\text{true,max}}\coloneqq\max\{K^{\text{true}}(1),\allowbreak K^{\text{true}}(0)\}$, in the sense that it results in the most conservative conclusions about ATE relative to any other prior consistent with our baseline assumptions about the potential outcomes. We establish this in appendix LABEL:sec:moreBayesian, alongside additional discussion of the Bayesian interpretation. In particular, we verify that $\underline{p}$ is a valid cdf on $\mathbb{R}_{+}$, and therefore a valid prior on $K^{\text{true,max}}$:
Proposition 1.
Suppose A1 holds. Then for any realization $\mathbf{X}$, $\underline{p}$ is monotonic, $\underline{p}$ is right continuous, $\lim_{K\rightarrow 0}\;\underline{p}(K)=0$, and $\lim_{K\rightarrow\infty}\;\underline{p}(K)=1$.
$K(\alpha)$ is therefore the smallest value of $K$ such that $K^{\text{true,max}}\leq K$ occurs with probability at least $1-\alpha$, according to the “robust” prior $\underline{p}$, which implies the same would be true for any other prior consistent with our baseline assumptions about the potential outcomes. So $[0,K(\alpha)]$ is a Bayesian credible set for $K^{\text{true,max}}$ with respect to $\underline{p}$; it is a set which contains $K^{\text{true,max}}$ with probability about $1-\alpha$ (the true probability is slightly higher than $1-\alpha$ since $\underline{p}$ is not continuous). And, again, this would also be true for any other prior consistent with our baseline assumptions. Consequently, $\Theta_{I}(K(\alpha))$ can be interpreted as a robust $(1-\alpha)$-credible set for the identified set for ATE. Since the identified set must contain the true ATE when the model is not falsified, this implies that $\Theta_{I}(K(\alpha))$ is also a robust $(1-\alpha)$-credible set for ATE itself.
Our approach is based on obtaining a distribution over identified sets, rather than a posterior distribution for ATE itself. This distinction is well known from the previous literature on Bayesian inference in partially identified settings (e.g., Kline and Tamer 2016 and the references therein). This analysis nonetheless allows us to make some probabilistic statements about ATE. For example, in figure 1, we can say that there is at least a 90% chance that ATE is in $\Theta_{I}(K(0.1))=[-6,22]$, and that there is at least a 65% chance that ATE is non-negative, both according to our randomization-based prior distribution $\underline{p}$.
Third, $\Theta_{I}(K(\alpha))$ has a frequentist interpretation as a design-based confidence set. To make the frequentist thought experiment explicit, write $\mathcal{C}(\mathbf{Y}(1)\times\mathbf{X}+\mathbf{Y}(0)\times(\mathbb{1}-\mathbf{X}),\mathbf{X})=\Theta_{I}(K(\alpha))$ to emphasize that both the identified set $\Theta_{I}(K)$ and $K(\alpha)$ depend on the vector of realized treatments $\mathbf{X}$ and the potential outcome vectors $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$. Let $\theta\big{(}\mathbf{Y}(1),\mathbf{Y}(0)\big{)}\coloneqq\frac{1}{N}\sum_{i=1}^{N}Y_{i}(1)-Y_{i}(0)$ denote the average treatment effect functional.
Theorem 3.
Suppose A1 and A3 hold. Then
$\inf_{\mathbf{Y}(1),\mathbf{Y}(0)\in[y_{\text{min}},y_{\text{max}}]^{N}}\;\mathbb{P}_{\text{design}}\Big{(}\mathcal{C}(\mathbf{Y}(1)\times\mathbf{X}^{\text{new}}+\mathbf{Y}(0)\times(\mathbb{1}-\mathbf{X}^{\text{new}}),\mathbf{X}^{\text{new}})\ni\theta\big{(}\mathbf{Y}(1),\mathbf{Y}(0)\big{)}\Big{)}\geq 1-\alpha.$
Keep in mind that $\mathbf{X}^{\text{new}}$ is the only random vector here. Theorem 3 shows that $\Theta_{I}(K(\alpha))$ is a $100(1-\alpha)$design-based confidence set. That is, across repeated random assignments of treatment, it will contain the true parameter with probability at least $1-\alpha$. In particular, this result holds for any fixed, finite population size $N$ that satisfies A1.
3.3 The value of randomization
In section 2.3 we showed that, in a finite population of any size, randomization does not have any identifying power because it does not guarantee level of ex-post balance. This result is overly pessimistic, however, because it does not take into account the impact that randomization has on our beliefs about balance. Our next result shows that we can obtain precise conclusions in large finite populations if we take these beliefs into account.
Theorem 4.
Consider a sequence of finite populations, $\{(Y_{i}(0)_{N},Y_{i}(1)_{N}):i=1,\ldots,N\}$. Assume that for each $x\in\{0,1\}$ there is a constant $\mu(x)\in\mathbb{R}$ such that $\frac{1}{N}\sum_{i=1}^{N}Y_{i}(x)_{N}\rightarrow\mu(x)$ as $N\rightarrow\infty$. Suppose A1 holds and that the bounds $y_{\text{min}}$ and $y_{\text{max}}$ do not depend on $N$. Assume $N_{1}/N\rightarrow\rho$ for some constant $\rho\in(0,1)$. Suppose that for each $N$, a vector of treatment assignments $\mathbf{X}_{N}$ is drawn from the random vector $\mathbf{X}_{N}^{\text{new}}$ that satisfies A3. Then:
  1. 1.
[…]
  2. 2.
    Let $\text{ATE}_{N}$ denote the finite population ATE. For any $\alpha\in(0,1)$,
[…]
Theorem 4 shows that a design-based sensitivity analysis will eventually yield very precise conclusions in a large population, whereas an analysis based solely on identification does not. Specifically, the first part of theorem 4 shows that all the mass of the worst case design-probability of the true magnitude of imbalance must eventually be to the left of the breakdown point. By definition of the breakdown point, this means that conclusions about the sign of ATE based on the exact balance assumption are guaranteed to be robust to failures of exact balance in sufficiently large populations.
The second part of theorem 4 shows that the sets $\Theta_{I}(K(\alpha))$ eventually collapse on the true ATE. Recall that $\Theta_{I}(K(\alpha))$ can be interpreted as a confidence set, by theorem 3. So this part of theorem 4 can be interpreted as saying that these confidence sets have statistical power—they are not trivially valid. Put differently, if you used $\Theta_{I}(K(\alpha))$ to construct a design-based test of the hypothesis $H_{0}:\text{ATE}_{N}=\theta_{0}$ then this test is consistent; it will reject false nulls with high design-probability when $N$ is large enough.
3.4 Measuring the strength of evidence: $\underline{p}(K^{\text{bp}})$ as an alternative to $p$-values
$p$-values are commonly interpreted as quantitative measures of the evidence against the null hypothesis, with small values interpreted as “strong evidence” against the null and larger values interpreted as weaker evidence against it. This is a controversial interpretation (e.g., see the entire 2019 special issue of The American Statistician on “a world beyond $p<0.05$”). In light of this debate about $p$-values, our measure $\underline{p}(K^{\text{bp}})$ can be viewed as an alternative summary statistic that has a well justified interpretation as a quantitative measure of how much evidence the data provides for a specific conclusion. In particular, because our analysis has a Bayesian interpretation, it allows us to make certain probabilistic statements about the true parameter itself. For example, when the naive difference in means estimate is positive, $\underline{p}(K^{\text{bp}})$ is a lower bound on the probability that the true ATE is nonnegative. Here “probability” refers to our randomization-based prior distribution.
This interpretation is particularly useful in applications with very small population sizes, where there can be a lot of uncertainty. In such settings, traditional hypothesis tests with conventional choices of the size $\alpha$ will often lead to a failure to reject the null, which is usually interpreted as “no evidence either way” given the lack of power arising from the small population size. However, our methods show that it is still possible to provide meaningful quantitative measures of how much evidence there is for conclusions about ATE even in small data sets. We give specific empirical examples in section 8.
3.5 Computing bounds on the design-probability of $K$-approximate balance
Our approach requires users to compute the function $\underline{p}(K)$, which involves solving an optimization problem over $N$ variables. In appendix LABEL:sec:MILPappendix, we show that $\underline{p}(K)$ is the optimized value in a mixed integer linear programming (MILP) problem. Consequently, standard software for solving these problems can be used. However, as we discuss in section 4, the MILP solver can often be very slow. So we also recommend that users try alternative solvers to compute $\underline{p}$. We have found that genetic algorithms (Kochenderfer and Wheeler 2019, pages 148–156) work exceptionally well in this setting, delivering results that are very close to the correct solution from MILP but in a small fraction of the time. See section 4 for a discussion of the numerical evidence.
A second issue concerns computation of the objective function, which involves a summation over the set of all possible treatment assignments. For small values like $N=10$ this is feasible but it can become infeasible for moderate values like $N=100$. This issue is well known in design-based inference; e.g., sec. 5.8 in Imbens and Rubin (2015). Like them, we address this issue by sampling from the set all treatment assignments and using this sample to approximate the objective function. In section 4 we show that our results are quite insensitive to the choice of sample size, so long as it is not too small.
4 Numerical Illustration
Next we use simulated data to illustrate the design-based sensitivity analysis of section 3. We generated a nested sequence of five populations, with $N\in\{10,20,40,100,400\}$. These populations have $[y_{\text{min}},y_{\text{max}}]=[0,1]$ and heterogeneous treatment effects such that $\text{ATE}_{N}$ converges to 0.25 as $N$ gets large. We assigned treatment via A3 such that $N_{1}=N/2$ for all $N$ while ensuring the populations are nested. In this illustration the assignment happens only once. Appendix LABEL:sec:numericalAppendix gives further details on how we generated the data.
[…]
(a) Convergence of $\underline{p}$ to a step function at zero as population size $N$ increases.
(b) Convergence of the bounds on ATE, $\Theta_{I}(K(\alpha))$, as population size $N$ increases. The lightest gray line is $N=10$ while the darkest line is $N=400$.
Figure 2: Illustration of the convergence results of theorem 4.
These results used a genetic algorithm (GA) to solve the optimization problem (3). Next we verify that this algorithm is obtaining accurate results by comparing its output with output from the mixed-integer linear programming (MILP) approach that we discussed in section 3.5. For $N=10$, the left plot in figure 3 shows the function $\underline{p}$ obtained using the genetic algorithm as a solid line, and the same function obtained using mixed-integer linear programming as a dashed line. The two lines are very similar, showing that the genetic algorithm is able to closely match the output of the MILP approach, despite being substantially faster. Specifically, for this plot the genetic algorithm took about 4 minutes, whereas MILP took about 34 hours. Similarly, the right plot in figure 3 shows $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$, as obtained by both algorithms. Again, the genetic algorithm is able to closely match the output from MILP. Finally, appendix figure LABEL:fig:robustness_batch_size shows that our results are robust to the choice of the number of treatment assignments used to approximate the objective function.
Figure 3: Assessing accuracy of the genetic algorithm: $\underline{p}$ vs $K$ (left plot), $\Theta_{I}(K(\alpha))$ vs $1-\alpha$ (right plot) for GA (solid line) and MILP (dashed line), $N=10$.
5 The Role of Balance in Covariates
Our analysis so far has ignored the role of covariates (sometimes called “attributes”). We discuss them next. For each unit $i\in\mathcal{I}$, let $W_{i}$ denote a vector of covariate values. Let $\mathbf{W}=(W_{1},\ldots,W_{N})$ denote the collection of covariate values for all units in the population. In practice, researchers commonly examine the observed magnitude of balance in covariates across the treatment and control groups. In this section we give a new formal justification for this kind of covariate balance analysis: Observed imbalances in covariates can be use to identify the unobserved realized imbalance in potential outcomes. Informally, this additional information arises when the covariates are predictive of potential outcomes. For this idea to have identifying power, however, we must make an explicit assumption on this relationship. There are many different formal assumptions one could consider. For brevity we focus on a particularly simple one here, but it would be useful to explore variations in future work. We illustrate this analysis empirically in section 8.2.
Let $\mathbf{Q}=(q(W_{1})^{\prime},\ldots,q(W_{N})^{\prime})^{\prime}$ be a $N\times\text{dim}(q(W_{i}))$ matrix of known transformations of the observed covariates. Assume $\mathbf{Q}^{\prime}\mathbf{Q}$ is invertible. Let $\beta(x)\coloneqq(\mathbf{Q}^{\prime}\mathbf{Q})^{-1}\mathbf{Q}^{\prime}\mathbf{Y}(x)$ denote the population regression coefficient from OLS of potential outcomes onto the transformed covariates. Let $U_{i}\coloneqq Y_{i}(x)-q(W_{i})^{\prime}\beta(x)$ denote the regression residual for unit $i$. Let $R_{Y(x)\sim q(W)}^{2}\coloneqq 1-\operatorname*{var}(U)/\operatorname*{var}(Y)$, where for any vector $(A_{1},\ldots,A_{N})$ we let $\overline{A}\coloneqq\frac{1}{N}\sum_{i=1}^{N}A_{i}$ denote its mean and $\operatorname*{var}(A)\coloneqq\frac{1}{N}\sum_{i=1}^{N}(A_{i}-\overline{A})^{2}$ denote its variance. Assume $q(W_{i})$ includes a constant, so that $\overline{U}=0$ by construction, since these are OLS residuals. Not all potential outcomes are observed. Hence $\beta(x)$, $(U_{1},\ldots,U_{N})$, and $R_{Y(x)\sim q(W)}^{2}$ are not point identified. Instead, we make assumptions about them, as follows.
Assumption A4 (Predictive covariates).
For each $x\in\{0,1\}$, $R_{Y(x)\sim q(W)}^{2}\geq\lambda$ for a known $\lambda\in[0,1]$.
$\lambda$ is a sensitivity parameter that controls the minimal predictive power of the covariates relative to unobserved variables that determine variation in potential outcomes. Larger values of $\lambda$ require a closer connection between potential outcomes and covariates. Since some potential outcomes are known, not all values of the sensitivity parameter $\lambda$ are consistent with the data. The largest value of $\lambda$ that is consistent with the data and all of the maintained assumptions is called the falsification point (Masten and Poirier 2021). Any $\lambda$ values below this point will lead to a non-falsified model, and therefore a non-empty identified set for ATE. Note that we could use different $\lambda$ values for each potential outcome; we use a common value for simplicity.
A4 has two implications: First, it has identifying power for potential outcomes $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$. Second, above and beyond its identifying power, it affects the worst case design probability of balance, $\underline{p}(K)$. That’s because it imposes a constraint on the unobserved values of potential outcomes. We consider each of these implications next.
5.1 The identifying power of covariates
In section 2.2 we derived an analytical expression for the finite population identified set for ATE under $K$-approximate mean balance assumption A2 alone. In this section we instead provide a numerical procedure for computing the identified set for ATE under the combined assumptions of bounded outcomes (A1), $K$-approximate mean balance (A2), and predictive covariates (A4).
Specifically, the upper bound on ATE solves
$\max_{\mathbf{Y}(1),\mathbf{Y}(0)\in[y_{\text{min}},y_{\text{max}}]^{2N}}\;\frac{1}{N}\sum_{i=1}^{N}\big{(}Y_{i}(1)-Y_{i}(0)\big{)}$
subject to (i) the data constraints that $Y_{i}(X_{i})=Y_{i}$ for all $i=1,\ldots,N$, (ii) the approximate $K$-means balance constraint
$-K\leq\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)\leq K,$
and (iii) the predictive covariates constraint $R_{Y(x)\sim q(W)}^{2}\geq\lambda$, which is equivalent to
$\sum_{i=1}^{N}\big{(}Y_{i}(x)-q(W_{i})^{\prime}\beta(x)\big{)}^{2}\leq N(1-\lambda)\operatorname*{var}(Y)$ (4)
where $\beta(x)=(\mathbf{Q}^{\prime}\mathbf{Q})^{-1}\mathbf{Q}^{\prime}\mathbf{Y}(x)$. The lower bound can be obtained by computing the minimum rather than the maximum. This is a quadratically constrained linear program since the objective function is linear, and all constraints are linear except (4), which is quadratic in the residuals. It is convex, which implies that it can be computed quickly and reliably in practice.
5.2 The impact of covariates on the worst case design probabilities of balance
The predictive covariates assumption also weakly increases the worst case probability that the $K$-approximate mean balance assumption A2 holds (so long as the model is not falsified), since it imposes additional constraints on the optimization problem. That is, we modify the definition of $\underline{p}$ to also impose the constraint in equation (4):
[…]
The additional constraint does not meaningfully affect the computational time.
5.3 Interpretation and discussion
Above we gave a new reason to study covariate balance: If observed covariates are linked to unobserved potential outcomes, then observed imbalances in covariates can inform us about realized imbalances in potential outcomes that occurred, even though treatment was randomly assigned. To build intuition, consider the case with a binary covariate $W_{i}$ and focus on the treated potential outcome. Let $q(W_{i})=(1,W_{i})^{\prime}$, and denote the corresponding components of $\beta(x)$ by $\beta_{0}(x)$ and $\beta_{1}(x)$. For any vector $(A_{1},\ldots,A_{N})$ let $\mathbb{E}(A\mid X=x)\coloneqq\left(\sum_{i=1}^{N}A_{i}\mathbbm{1}(X_{i}=x)\right)/\sum_{i=1}^{N}\mathbbm{1}(X_{i}=x)$. By the definition of the residuals $U_{i}(1)$, $\mathbb{E}[Y(1)\mid X=x]=\beta_{0}(1)+\beta_{1}(1)\mathbb{E}(W\mid X=x)+\mathbb{E}[U(1)\mid X=x]$ for $x\in\{0,1\}$ and hence
$\mathbb{E}[Y(1)\mid X=0]-\mathbb{E}[Y(1)\mid X=1]$
$=\beta_{1}(1)\big{(}\mathbb{E}(W\mid X=0)-\mathbb{E}(W\mid X=1)\big{)}+\big{(}\mathbb{E}[U(1)\mid X=0]-\mathbb{E}[U(1)\mid X=1]\big{)}.$
This equation shows that the magnitude of imbalance in potential outcomes depends directly on the magnitude of imbalance in covariates. It also depends on the magnitude of imbalance in the residuals, which is controlled by $\lambda$. In particular, for $\delta\coloneqq\sqrt{N(1-\lambda)\operatorname*{var}(Y)}$, $\mathbb{E}[U(1)\mid X=x]\in[-\delta,\delta]$ for each $x\in\{0,1\}$. Hence
$\big{|}\mathbb{E}[U(1)\mid X=0]-\mathbb{E}[U(1)\mid X=1]\big{|}\leq 2\delta.$
This bound combined with the observed imbalance in covariates directly constrain the imbalance in potential outcomes, which is the source of identifying power behind the covariates. A similar analysis applies to balance in $Y(0)$. Note that this discussion is for intuition only; section 5.1 describes how we obtain identified sets in general.
In section 3.3 we discussed how randomization guarantees approximate balance asymptotically. In contrast, randomization does not imply anything about the value of $R_{Y(x)\sim q(W)}^{2}$; it is a population level parameter that does not depend on how treatment is assigned. Consequently, we cannot use randomization to calibrate the magnitude of its lower bound, $\lambda$. However, keep in mind that assumptions like A4 are not necessary for drawing tight conclusions about ATE asymptotically; this is shown in the second part of theorem 4. Their main purpose is to provide tighter bounds in finite populations.
We conclude this section with several remarks about the literature. First, covariate balance is sometimes examined to test whether treatment was actually randomized. Here we simply assume treatment was in fact randomized. Second, the traditional design-based inference framework primarily uses covariates to motivate different choices of test statistics, with the goal of increasing the power of hypothesis tests. For example, see Imbens and Rubin (2015, section 5.9) or Zhao and Ding (2021). This approach, like ours, uses covariates to derive stronger conclusions about the parameter of interest. Mathematically, however, our approach uses covariates for identification and does not rely on hypothesis testing theory. Finally, covariates are also used to define the parameter of interest (e.g., Abadie et al. 2020). In principle this can be done in our approach too, but we leave this to future work (also see section 7.3).
6 Noncompliance
We have focused on the classical case where all units comply with their treatment assignment. In this section we extend the analysis to a finite population version of the Imbens and Angrist (1994) model (as in section 3 of Hong et al. 2020), allowing us to study settings with noncompliance. In the baseline case of exact balance, we show that the Wald estimand point identifies a realized local average effect of treatment on the treated (LATT) parameter. We then study finite population identification under $K$-approximate mean balance type assumptions. In the special case of one-sided noncompliance, we show that the finite population identified set for the realized LATT has a simple form that is analogous to classical large population results. We then show how to use this result to do design-based sensitivity analysis. In all of this analysis we allow for heterogeneous treatment effects (in contrast to some prior work on design-based instrumental variable analysis, such as Rosenbaum 1996).
6.1 Setup
As before, there is a finite population of units $i=1,\ldots,N$. For each unit $i$: Let $Y_{i}(1)$ and $Y_{i}(0)$ denote potential outcomes, $X_{i}(1)$ and $X_{i}(0)$ the binary potential treatments, $Z_{i}$ the realized value of a binary instrument (assigned treatment in the noncompliance setting), $X_{i}=X_{i}(Z_{i})$ the realized treatment, and $Y_{i}=Y_{i}(X_{i})$ the realized outcome. Note that we impose the exclusion restriction implicitly here and maintain it throughout this section. Without loss of generality, write $Y_{i}(x)=\beta_{i}\cdot x+U_{i}$ for $x\in\{0,1\}$, where we defined $U_{i}\coloneqq Y_{i}(0)$ and $\beta_{i}\coloneqq Y_{i}(1)-Y_{i}(0)$. Let $N_{1}=\sum_{i=1}^{N}\mathbbm{1}(Z_{i}=1)$ denote the number of units whose instrument value equals 1. We’ll call this the instrument-on group. Let $N_{0}=N-N_{1}$. We’ll call the set of units with $Z_{i}=0$ the instrument-off group. Let $\overline{Y}_{z=1}\coloneqq\frac{1}{N_{1}}\sum_{i:Z_{i}=1}Y_{i}$ denote the average outcome in the instrument-on group. Define $\overline{Y}_{z=0}$, $\overline{X}_{z=1}$, $\overline{X}_{z=0}$, $\overline{U}_{z=1}$, and $\overline{U}_{z=0}$ similarly.
6.2 Baseline identification
Define the compliance type variable
$T_{i}=\begin{cases}c&\text{if $X_{i}(1)=1,X_{i}(0)=0$}\\ a&\text{if $X_{i}(1)=1,X_{i}(0)=1$}\\ n&\text{if $X_{i}(1)=0,X_{i}(0)=0$}\\ d&\text{if $X_{i}(1)=0,X_{i}(0)=1$}.\end{cases}$
We maintain the following finite population version of the monotonicity / no defiers assumption in Imbens and Angrist (1994).
Assumption B1 (No defiers).
$T_{i}\neq d$ for all $i=1,\ldots,N$.
Similarly, we assume the following finite population version of relevance holds for the specific realization of treatment assignment that is observed.
Assumption B2 (Relevance).
$\overline{X}_{z=1}\neq\overline{X}_{z=0}$.
Let $\overline{T}_{1}(a)\coloneqq\frac{\sum_{i=1}^{N}\mathbbm{1}(Z_{i}=1)\mathbbm{1}(T_{i}=a)}{\sum_{i=1}^{N}\mathbbm{1}(Z_{i}=1)}$ denote the proportion of always takers in the instrument-on group. Likewise, let $\overline{T}_{0}(a)$ denote the proportion of always takers in the instrument-off group, and $\overline{T}_{1}(c)$ the proportion of compliers in the instrument-on group. Let $\overline{\beta}_{1}(a)=\left(\sum_{i:Z_{i}=1,T_{i}=a}\beta_{i}\right)/\allowbreak\sum_{i=1}^{N}\mathbbm{1}(Z_{i}=1)\mathbbm{1}(T_{i}=a)$ denote the average treatment effect among the instrument-on always takers. Define $\overline{\beta}_{0}(a)$ and $\overline{\beta}_{1}(c)$ similarly.
Lemma 1.
Suppose B1 (no defiers) and B2 (relevance) hold. Then
$\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}=\frac{\overline{U}_{z=1}-\overline{U}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}+\frac{\overline{T}_{1}(a)\overline{\beta}_{1}(a)-\overline{T}_{0}(a)\overline{\beta}_{0}(a)}{\overline{X}_{z=1}-\overline{X}_{z=0}}+\frac{\overline{T}_{1}(c)}{\big{(}\overline{T}_{1}(a)-\overline{T}_{0}(a)\big{)}+\overline{T}_{1}(c)}\;\text{\footnotesize$\overline{\beta}_{1}(c)$}.$
The left hand side of equation (6) is the finite population version of the Wald estimand. So this lemma decomposes the Wald estimand into three pieces that each depend on the magnitude of various realized imbalances. For our baseline analysis, consider the following exact mean balance assumption.
Assumption B3 (Exact balance).
(i) $\overline{U}_{z=1}=\overline{U}_{z=0}$, (ii) $\overline{T}_{1}(a)=\overline{T}_{0}(a)$, and (iii) $\overline{\beta}_{1}(a)=\overline{\beta}_{0}(a)$.
[…]
Proposition 2.
Suppose B1 (no defiers), B2 (relevance), and B3 (exact balance) hold. Then $\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}=\overline{\beta}_{1}(c)$.
Proposition 2 shows that $\overline{\beta}_{1}(c)$ is point identified in finite populations under exact balance. In particular, it equals the finite population Wald estimand. This result is a finite population version of the Imbens and Angrist (1994) result, with one slight difference: The point identified parameter $\overline{\beta}_{1}(c)\coloneqq\left(\sum_{i=1}^{N}\beta_{i}\cdot\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)$ is a realized local average effect of treatment on the treated (LATT)—it is the average unit level causal effect among treated compliers (recall that $Z_{i}=X_{i}$ for compliers). In contrast, the population LATE is $\overline{\beta}(c)\coloneqq\left(\sum_{i=1}^{N}\beta_{i}\mathbbm{1}(T_{i}=c)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)$. The LATT is an ex ante random parameter, since the set of units who will be treated, and thus who appear in the parameter’s definition, depends on the realization of treatment assignment. This is analogous to Rosenbaum’s (2001) finite population analysis of ATT, where he noted that the ATT parameter is also ex ante random (he called the ATT the “attributable effect”). Another slight difference from the standard super-population analysis is that when there is one-sided noncompliance (so there are no always takers), the ATT and LATT are the same, but they do not equal LATE because there can be non-treated compliers. Finally, note that we can decompose LATE as $\overline{\beta}(c)=p_{1}(c)\overline{\beta}_{1}(c)+p_{0}(c)\overline{\beta}_{0}(c)$ where $p_{1}(c)\coloneqq\left(\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)$ is the proportion of units who are treated, among all compliers, and likewise for $p_{0}(c)$. So if we further assume that the average treatment effect for compliers in the instrument-on group is the same as for compliers in the instrument-off group—$\overline{\beta}_{1}(c)=\overline{\beta}_{0}(c)$—then the finite population Wald estimand equals LATE. This additional condition is analogous to part (iii) of B3.
6.3 Design-based sensitivity analysis
Proposition 2 shows that the Wald estimand equals LATT under an exact balance assumption. However, as discussed in section 2, exact balance generally does not hold in small finite populations. Instead, we can apply the same ideas from earlier to perform a design-based sensitivity analysis: We can derive identified sets for LATT under approximate balance assumptions and then use randomization to calibrate the sensitivity parameters. Here we briefly sketch the analysis in the one-sided noncompliance case, which is particularly straightforward.
Assumption B4 (One-sided noncompliance).
$T_{i}\neq a$ for all $i=1,\ldots,N$.
Without always takers, the second term in equation (6) disappears, and the third term becomes $\overline{\beta}_{1}(c)$. Hence the only remaining term is a difference in non-treated potential outcomes, which we bound via the following assumption.
Assumption B5 ($K$-approximate mean balance for $Y(0)$).
There is a known $K\geq 0$ such that $|\overline{U}_{z=1}-\overline{U}_{z=0}|\leq K$.
The next result bounds the realized LATT as a function of $K$. Here we let $\pi\coloneqq\overline{X}_{z=1}-\overline{X}_{z=0}$ denote the first stage difference in means.
Theorem 5.
Suppose B1 (no defiers), B2 (relevance), B4 (one-sided noncompliance), and B5 ($K$-approximate mean balance for $Y(0)$) hold. Then the finite population identified set for $\overline{\beta}_{1}(c)$ is
$\left[\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}-\frac{K}{\pi},\;\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}+\frac{K}{\pi}\right].$
The identified set in theorem 5 is analogous to the classical large population identified sets where instrument exogeneity is relaxed at the population level. For example, see Bound et al. (1995) or Conley et al. (2012). The identified set in theorem 5 can be further adjusted to impose the bounded outcome assumption A1, like in theorem 1. Then, given an assumption on the design distribution of the instrument, we can construct a function similar to $\underline{p}$ in equation (3) and use this to calibrate the value of $K$. We omit the details for brevity. The general two-sided noncompliance case is more complicated, since it involves more than just a single balance condition (i.e., relaxations of the three conditions in B3). We conjecture that our analysis extends to this case but leave a full exploration to future work.
7 Extensions
7.1 Sampling
From section 2.2 onward we assumed for simplicity that there was no sampling—all units in the population are observed, and the only uncertainty is about the unknown potential outcomes. Here we briefly discuss two extensions: An analysis of sampling by itself, and an analysis the combines both sampling and random assignment (e.g., as in Abadie et al. 2020). In both cases we reframe the classical question of inferring population quantities from sample data as an identification problem, which allows us to provide a single approach to quantify uncertainty from missing units and from missing potential outcomes. Keep in mind that this extension will generally make bounds wider, since it accounts for an additional source of uncertainty.
First consider the sampling setup from section 1. The population are the numbers $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$. We only observe $n<N$ of these units, and the goal is to learn about the population mean $\overline{Y}\coloneqq\frac{1}{N}\sum_{i=1}^{N}Y_{i}$. Let $S_{i}\in\{0,1\}$ denote whether unit $i$ is sampled or not. From an identification perspective, sampling is a missing data problem—we observe the values $Y_{i}$ when $S_{i}=1$ but have no data whatsoever on values $Y_{i}$ with $S_{i}=0$. Consequently, if all we know are that all outcomes $Y_{i}$ lie in known bounds (an assumption similar to A1), then all we can say for sure about the population mean is that it lies in domain-of-consensus bounds analogous to $\Theta_{I}(\infty)$; this is the motivation behind including sample indicators in definition 1. However, we can shrink the identified set by making assumptions like
$\left|\frac{1}{n}\sum_{i=1}^{N}S_{i}Y_{i}-\frac{1}{N-n}\sum_{i=1}^{N}(1-S_{i})Y_{i}\right|\leq K,$
which is analogous to the $K$-approximate mean balance assumption A2. Under this assumption, we can derive identified sets for $\overline{Y}$ as a function of the sensitivity parameter $K$. Finally, suppose we know the sample was obtained via simple random sampling (SRS), for example. This is the sampling analog of uniform randomization (A3)—all possible samples of size $n$ from $N$ have equal probability of being selected. Then we can compute worst case design probabilities of imbalance, which we can use to calibrate the sensitivity parameter $K$. This allows us to perform a design-based sensitivity analysis for sampling.
Next consider the randomized experiment setting we focused on in this paper. Suppose that, in addition to random assignment, we only observe outcomes $Y_{i}$ for a sample of $n<N$ units. In this case we need to address the identification problem that arises from missing potential outcomes as well as missing data on some units altogether. This setting does not require any new conceptual ideas, and so we only discuss it briefly. Consider the average treated potential outcome. Let $n_{1}<n$ denote the number of treated units. Assume both $n$ and $n_{1}$ are fixed a priori, and both sampling and randomization are performed independently according to simple random sampling and uniform randomization. We observe the average treated potential outcome among sampled units who are treated, $\frac{1}{n_{1}}\sum_{i:S_{i}=1,X_{i}=1}Y_{i}(1)$. We do not know the average $\frac{1}{N-n_{1}}\sum_{i:S_{i}=0\text{ or }X_{i}=0}Y_{i}(1)$. However, we can make a $K$-approximate mean balance assumption that says this unobserved mean is not too far from the observed one. Then we can use our sampling and randomization assumptions to calibrate the value $K$. The same analysis can be done for the non-treated potential outcome, and they can be combined to do a design-based sensitivity analysis for the population ATE.
Our approach treats sampling and randomization symmetrically—both lead to missing data which can be thought of as a finite population identification problem, and therefore analyzed using the tools of partial identification. However, there is a distinction between the two, which helps reconcile the difference between our definition 1 and definitions like Koopmans’ (1949) which assume away sampling uncertainty: Let $\mathcal{D}$ be the set of all designs, which are mappings from $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ into joint probabilities over sampling and treatment assignments, $(\mathbf{S},\mathbf{X})$. Say a parameter is point identifiable with respect to a set of assumptions $\mathcal{P}$ if there exists a design such that, for any $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ satisfying the assumptions $\mathcal{P}$, and for any realization $(\mathbf{S},\mathbf{X})$ from that design, the subsequent identified set $\Theta_{I}$ is a singleton. Then the population mean $\overline{Y}$ is point identifiable with respect to A1, since we can in principle sample all units with probability 1. But in any fixed dataset with $n<N$ it will only be partially identified. In contrast, ATE is not point identifiable with respect to A1 since we can never treat everyone and treat no-one.
7.2 Variations on the bounded outcomes assumption
Since averages are influenced by outlier units, obtaining nontrivial bounds on ATE usually requires some kind of assumption that restricts the magnitude of outlier values. Throughout this paper we used a simple uniform bound $[y_{\text{min}},y_{\text{max}}]$ on potential outcomes, assumption A1. First, it is important to recognize that this is a substantive identifying assumption, not a regularity condition—if these bounds are large then the researcher is explicitly allowing for large outliers and hence potentially large magnitudes of imbalance. Researchers who do not want to allow for such outliers must explicitly rule them out by assumption.
Second, the specific form of the assumption we used can be replaced or augmented with a variety of similar assumptions, and all of our methods will continue to apply. Here we give just a few examples: (i) A simple extension is to allow unit specific bounds, $Y_{i}(x)\in[y_{\text{min},i},y_{\text{max},i}]$. We use this version in one of our empirical applications, where the units are groups and the outcome depends on group size. (ii) One could impose a bounded unit level treatment effect assumption: $|Y_{i}(1)-Y_{i}(0)|\leq M$ for all $i\in\mathcal{I}$, where $M$ is a known sensitivity parameter. This assumption implies unit-specific bounds on the unobserved potential outcomes: $Y_{i}(x)\in[Y_{i}-M,Y_{i}+M]$. (iii) Alternatively, one could assume the sum of unit level treatment effect magnitudes is bounded: $\sum_{i=1}^{N}|Y_{i}(1)-Y_{i}(0)|\leq M$ for a known $M$. This would allow for some units to have very large treatment effects, so long as not too many do. This condition implies that ATE can be no larger than $M/N$. (iv) Or one could restrict the population variance of potential outcomes to be no larger than a known $M$. This assumption implies that $Y_{i}(x)$ is in the interval $\frac{1}{N_{x}}\sum_{i=1}^{N}Y_{i}\mathbbm{1}(X_{i}=x)\pm 2\sqrt{N_{x}M}$. Hence it does not require users to specify a priori bounds $[y_{\text{min}},y_{\text{max}}]$; they only must choose $M$. (v) Finally, the covariate restrictions in section 5 can also be viewed as one way of restricting the magnitude of outliers.
7.3 Distributional balance and parameters beyond ATE
Our analysis has defined balance based on differences in means. There are many other ways to measure the balance in a variable across the treatment and control groups; see chapter 14 of Imbens and Rubin (2015). For example, letting $\mathbb{P}^{\text{true}}(Y(x)\leq y\mid X=x)\coloneqq\frac{1}{N_{x}}\sum_{i:X_{i}=x}\mathbbm{1}[Y_{i}(x)\leq y]$ (see appendix LABEL:sec:popDistributions), we could consider the assumption
$\sup_{y\in\mathbb{R}}\big{|}|\mathbb{P}^{\text{true}}(Y(x)\leq y\mid X=1)-\mathbb{P}^{\text{true}}(Y(x)\leq y\mid X=0)\big{|}\leq K$ (7)
which bounds the sup-norm distance between the population distributions of potential outcomes in the treatment and control groups. We could then derive identified sets for the parameter of interest under this assumption, for a fixed $K$. Given a randomization design, we could them compute the worst case design-probability that (7) holds, which would lead to a design-based sensitivity analysis. This alternative approach suggests that analyzing different forms of balance might lead to tighter bounds than those based on mean balance alone. Thus the choice of balance metric in our analysis could be thought of as analogous to the choice of the test statistic in classical randomization tests. Whether such gains are possible will likely depend on the parameter of interest; for example, it is not clear if such distributional notations of balance have any additional implications for mean parameters like ATE. Alternative balance metrics may be more appropriate for studying parameters beyond ATE, however. For example, the sup-norm distance in equation (7) will likely work well for identification of quantile treatment effects (QTEs) since these are defined using inverses of the unconditional population cdfs $\mathbb{P}^{\text{true}}(Y(x)\leq y)\coloneqq\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[Y_{i}(x)\leq y]$. We leave a full exploration of these questions to future work.
8 Empirical Applications
In this section we illustrate our approach in two empirical applications with particularly small populations ($N=10$ and $N=17$). While our methods apply to any population size, and are feasible for larger population sizes (see appendix LABEL:sec:LaLonde for a third application with $N=722$), these two applications show that it is still possible to do meaningful inference in small datasets. In appendix LABEL:sec:AlternativeEmpiricalCIs we show three alternative frequentist confidence intervals for comparison.
8.1 Getting parents to pick their kids up on time
Our first application uses data from Gneezy and Rustichini (2000, Journal of Legal Studies). This paper is well known, with about 3400 Google Scholar citations as of June 2025, and has been frequently discussed in popular press books like Freakonomics. They study day care centers in Israel, where administrators were frustrated with parents showing up late to pick up their kids. They asked: Would a monetary fine incentivize parents to be on time? The population is 10 centers. The treatment is the introduction of a center-wide late fee. 6 centers were treated and 4 were not. The outcome variable is the number of late parents in a week. The logical lower bound is zero and the logical upper bound is five times the number of kids in that center (assuming every parent is late every day of the week). We assume that regardless of treatment, on average, each child has a late parent at most once per week. That is, we set $y_{\text{max},i}$ equal to the number of children in center $i$ (see (i) in section 7.2). In the data there are between 28 and 37 kids per center. We let $y_{\text{min},i}=0$ for all units.
The authors gathered baseline data on outcomes for 4 weeks. The fee was introduced at treated centers at the beginning of week 5. It was removed at the beginning of week 17. The authors gathered another 4 weeks of data after removal of the fee, for a total of 20 weeks of data. Their table 1 provides the full dataset. For simplicity we only use data from one post-treatment week, week 19. It would be interesting to study how to extend our results to use the time series data as well, but we leave this to future work.
The ATE point estimate is $13.7$ late parents, suggesting that the addition of a fee increased the number of late parents per week by about 14. To quantify the uncertainty around this estimate, we conduct a design-based sensitivity analysis. Figure 1 in the introduction shows the results, which we briefly discussed already. First consider the breakdown point, $\underline{p}(K^{\text{bp}})=65$%. So there is at least a 65% chance that the ATE is non-negative, according to our randomization based prior distribution. Although this does not attain conventional levels of “confidence”, like 95%, it is nonetheless a non-trivial inference given that our dataset only contains 10 units in it. This conclusion can also be seen in figure 1(b), which plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$. All sets with probabilities smaller than 64% contain only positive values. Moreover, even for large probabilities, the sets $\Theta_{I}(K(\alpha))$ are still mostly in the positive region. For example, the 90% set is about $[-6,22]$. There is at least a 90% chance that the true ATE is in this set. Overall, these findings suggest that there is reasonably strong evidence that the ATE is in fact positive.
8.2 The long run adoption of management practices
Our second application uses data from Bloom et al. (2013, QJE) and Bloom et al. (2020, AEJ: Applied). These are influential papers with about 2400 total Google scholar citations as of June 2025. These papers asked whether large observed differences in productivity across firms are driven by variation in firms’ management practices. To answer this, they ran a randomized experiment in a population of 17 woven cotton fabric firms in India. These were large and old firms, with an average of 270 employees per firm and an average age of 20 years old at baseline. Their control group received a one-month diagnostic about their management practices. The treatment group received the diagnostic plus four months of support for implementing the management changes. They consider many different outcomes of interest, but we will focus on the long run outcome from their 2020 paper. Specifically, the outcome variable is the proportion of 38 management practices adopted in 2017, which was about 8 to 9 years after they received treatment. Since this outcome is a proportion, we set $[y_{\text{min}},y_{\text{max}}]=[0,1]$.
(a) Top: $\Theta_{I}(K)$, bounds on ATE as a function of the maximum magnitude of ex post imbalance, $K$. Bottom: $\underline{p}(K)$.
(b) $\Theta_{I}(K(\alpha))$, bounds on the ATE as a function of $1-\alpha$. The darkest line is $\lambda=0$, the lightest line is $\lambda=0.4$, and the line in the middle is $\lambda=0.2$.
Figure 4: Quantifying the uncertainty in the impact of management interventions on long run adoption of management practices.
Some of the 17 firms in the population have multiple plants. Treatment was assigned and administered at the plant-level, with at most 1 plant per firm treated. This implies that there are non-treated plants at firms with a different treated plant. The authors use this data to study within-firm spillovers. For simplicity we ignore all data from non-treated plants at treated-firms. This gives us a dataset where each unit is a single plant. There are 11 treated plants and 6 control plants.
The ATE point estimate is 0.13, suggesting that providing 4 months of support for changing management practices leads to a 13 percentage point increase in the proportion of practices that are still in place about one decade later. The authors emphasize that it is important to also measure the uncertainty associated with this point estimate, however:
“The major challenge of our experiment was its small cross-sectional sample size. We have data on only 28 plants across 17 firms. To address concerns over statistical inference in small samples, we implemented permutation tests whose properties are independent of sample size.” (page 4, Bloom et al. 2013)
That is, to deal with the small population size they performed exact tests of the sharp null of no unit level treatment effects, assuming uniform randomization. In the 2013 paper they also used time series methods that are valid for fixed $N$ and large $T$, which we ignore since they cannot be used with the long run outcomes data. To complement their original results, we implement our design-based sensitivity analysis. Figure 4 shows the main results. First consider the breakdown point, $\underline{p}(K^{\text{bp}})=20$%. So there is at least a 20% chance that the ATE is non-negative, according to our randomization based prior distribution. So with this dataset it is unlikely that potential outcomes would be balanced enough to ensure that $K$ is small enough that we can rule out negative ATE values. This is also shown in the outer-most bounds of figure 4(b), which plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$. Here we see that the sets all contain negative numbers for probabilities larger than 20%. For example, the 90% interval is $[-0.28,0.49]$. So there is at least a 90% chance that ATE is inside this interval, because this interval can be interpreted as a Bayesian credible set.
We can obtain tighter bounds by adding assumptions about covariates, as in section 5. Figure 4(b) also shows bounds on ATE that impose the predictive covariates restriction A4 for $\lambda=0$ (no covariate restrictions), $0.2$, and $0.4$, going from the darkest, outer-most bands to the lightest, inner-most bands. We stop at $\lambda=0.4$ since larger values are generally falsified. Following the authors’ analysis (table 2 in Bloom et al. 2020), we only use a single covariate, baseline management score in 2008, which is essentially a lagged outcome, and specify a simple linear model $q(W_{i})=(1,W_{i})^{\prime}$. For this choice, the predictive covariates assumption tightens the bounds somewhat, bringing the breakdown point up to about 45%, compared to its unconstrained value of 20%. The bounds could potentially be further tightened by using additional covariates (e.g., those in their table 1), or by making additional assumptions directly on the unobserved potential outcomes, as in section 7.2.
9 Conclusion
In this paper we studied identification in finite populations. We formally showed that, for any finite population size, randomization has no identifying power, because it does not guarantee any level of balance in unobservables. Nonetheless, we showed how to use randomization to derive empirical objective beliefs about the ex post level of balance. By combining finite population identified sets with the empirical objective worst case prior, we showed how to conduct a design-based sensitivity analysis, a new approach to quantifying uncertainty in finite populations. This approach yields uncertainty intervals that have (1) identification, (2) Bayesian, and (3) frequentist interpretations. From the frequentist perspective, our confidence intervals allow for arbitrary heterogeneous treatment effects, but do not rely on asymptotics. And we showed that our confidence intervals are consistent; they have statistical power. The key idea is that we are not doing inference based on hypothesis test inversion. Instead, we combine partial identification analysis with ideas from design-based inference to construct uncertainty intervals. Our approach also gave a new motivation for examining covariate balance, which can be used to shrink identified sets. We applied these ideas in both (1) classical randomized experiments and (2) the instrumental variable setting, focusing on one-sided noncompliance.
Many open questions remain: In section 7 we sketched extensions to sampling, distributional measures of balance, and parameters beyond ATE, but left a full analysis for future work. We have focused on uniform randomization throughout this paper, but our approach likely generalizes to other types of randomization designs. One can also ask what optimal experimental designs are, given that the analyst will perform a design-based sensitivity analysis. Finally, we have focused on relatively simple settings with randomization by design, but an important next step is to extend this analysis to other settings like difference-in-differences or synthetic controls.
Appendix A Proofs of main results
Proof of theorem 1.
Here we use the notation of appendix LABEL:sec:popDistributions. By iterated expectations, $\mathbb{E}[Y(x)]=\mathbb{E}(Y\mid X=x)\mathbb{P}^{\text{data}}(X=x)+\mathbb{E}[Y(x)\mid X=1-x]\mathbb{P}^{\text{data}}(X=1-x)$. By A1 and A2, $\mathbb{E}[Y(x)\mid X=1-x]\in[\max\{y_{\text{min}},\mathbb{E}(Y\mid X=x)-K\},\min\{y_{\text{max}},\mathbb{E}(Y\mid X=x)+K\}]$. Substitute to get the bounds in the theorem. Sharpness follows as in Manski (1990), with the minor addition that knowledge of unit identifiers in $\mathbf{P}^{\text{data}}$ has no identifying power. Specifically, any values of $\mathbb{E}[Y(1)\mid X=0]\in[y_{\text{min}},y_{\text{max}}]$ and $\mathbb{E}[Y(0)\mid X=1]\in[y_{\text{min}},y_{\text{max}}]$ are consistent with $\mathbf{P}^{\text{data}}$ because these averages depend solely on values of potential outcomes that are not present in the data. ∎
Proof of theorem 2.
From the proof of theorem 1 we know that for any element $\theta\in\Theta_{I}(\infty)$, there exists a population matrix $\mathbf{P}$ that is consistent with A1 and $\mathbf{P}^{\text{data}}$ and has $\theta=\theta(\mathbf{P})$. We now also know, however, that $\mathbf{X}$ is a realization from the distribution $\mathbb{P}_{\text{design}}$. However, this knowledge does not constrain the unobserved values of potential outcomes in any way, since $\mathbb{P}_{\text{design}}$ does not depend on potential outcomes. Hence it is still consistent with $\mathbf{P}$ being the true value of the population matrix. ∎
Recall that $K^{\text{true}}(x)$ (eq (2)) denotes the true magnitude of imbalance in the $x$-potential outcome. Let $K^{\text{true,new}}(x)$ denote the same expression but with the random variables $X_{i}^{\text{new}}$ replacing $X_{i}$ for all $i\in\mathcal{I}$.
Proof of proposition 1.
$p(\cdot,\mathbf{Y}(1),\mathbf{Y}(0))$ is the cdf for the random variable $\max\{\allowbreak K^{\text{true,new}}(1),\allowbreak K^{\text{true,new}}(0)\}$, and hence has all the properties of a cdf. We will show that these properties are preserved once we take the infimum over $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$.
1. Monotonicity: Since $p(\cdot,\mathbf{Y}(1),\mathbf{Y}(0))$ is a cdf, it is monotonic: $p(K_{1},\mathbf{Y}(1),\mathbf{Y}(0))\leq p(K_{2},\mathbf{Y}(1),\mathbf{Y}(0))$ for any $K_{1}\leq K_{2}$, for all $\mathbf{Y}(1),\mathbf{Y}(0)$. Taking the infimum preserves the inequality, to get $\underline{p}(K_{1})\leq\underline{p}(K_{2})$ for any realization $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$.
2. $\underline{p}$ is right continuous: $p(\cdot,\mathbf{Y}(1),\mathbf{Y}(0))$ is monotonic and right continuous, which implies that it is upper semicontinuous everywhere. Moreover, the pointwise infimum of an upper semicontinuous function is still upper semicontinuous. Thus $\underline{p}(\cdot)$ is upper semicontinuous. And from part 1, $\underline{p}(\cdot)$ is monotonically increasing. Now apply lemma LABEL:lemma:upperSemi.
3. Limits: Convergence to 1 as $K\rightarrow\infty$: Let $K\geq y_{\text{max}}-y_{\text{min}}$. Then A1 implies that for all dgps $\mathbf{Y}(1),\mathbf{Y}(0)$, for all realizations $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $|\overline{Y}_{1}(x)-\overline{Y}_{0}(x)|\leq K$, where $\overline{Y}_{g}(x)\coloneqq\frac{1}{N_{g}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=g)$ for $g\in\{0,1\}$. Thus $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=1$ for all such $K$. Since this holds for all $\mathbf{Y}(1),\mathbf{Y}(0)$, taking the infimum does not change the result.
Convergence to 0 as $K\rightarrow 0$: We want to show that, for each realization $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $\lim_{K\rightarrow 0}\underline{p}(K)=0$. We’ll show something stronger: There is an $\varepsilon>0$ such that $\underline{p}(K)=0$ for $K\in[0,\varepsilon)$. Recall that $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=\mathbb{P}(K^{\text{true,new}}(1)\leq K,K^{\text{true,new}}(0)\leq K)$. And $\underline{p}$ is the infimum of this term over all $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ satisfying A1. Thus it suffices to find (i) a single dgp $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ that is consistent with the data and (ii) a small enough value of $K$ such that, for that dgp, just one of the two conditions $K^{\text{true,new}}(x)\leq K$ cannot hold. Consider the non-treated potential outcomes. Order the observations so the first $N_{1}$ units are treated. We observe $\mathbf{Y}(0)_{1:N_{1}}$ in the data. It suffices to find values $\mathbf{Y}(0)_{N_{1}+1:N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ and an $\varepsilon>0$ such that $|(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{1:N_{1}})/N_{1}-(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{N_{1}+1:N})/N_{0}|>\varepsilon$, where $\boldsymbol{\iota}$ are vectors of 1’s. This can be done by choosing all components of $\mathbf{Y}(0)_{N_{1}+1:N}$ equal to $y_{\text{min}}$ or all components equal to $y_{\text{max}}$; note that in each case the average of the unobserved $Y_{i}(0)$’s is also equal to either $y_{\text{min}}$ or $y_{\text{max}}$. And note that $c\coloneqq(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{1:N_{1}})/N_{1}$ is a constant that is fixed in the data, with $c\in[y_{\text{min}},y_{\text{max}}]$. If $c=y_{\text{min}}$ then pick the unobserved $Y_{i}(0)$’s all equal to $y_{\text{max}}$ while if $c=y_{\text{max}}$ we pick the unobserved $Y_{i}(0)$’s all equal to $y_{\text{min}}$. In either case we set $\varepsilon=y_{\text{max}}-y_{\text{min}}>0$. If $c$ is strictly inside $[y_{\text{min}},y_{\text{max}}]$ then either choice for the $Y_{i}(0)$’s works; suppose we set them all to $y_{\text{min}}$. Then set $\varepsilon=|c-y_{\text{min}}|>0$. ∎
Proof of theorem 3.
For brevity, let $\mathcal{C}=\mathcal{C}(\mathbf{Y}(1)\times\mathbf{X}^{\text{new}}+\mathbf{Y}(0)\times(\mathbb{1}-\mathbf{X}^{\text{new}}),\mathbf{X}^{\text{new}})$ denote the random confidence set. Recall its realizations are sets $\Theta_{I}(K(\alpha))$. First note that
$\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K(\alpha)\text{ for $x\in\{0,1\}$})\leq\mathbb{P}_{\text{design}}\big{(}\mathcal{C}\ni\theta(\mathbf{Y}(1),\mathbf{Y}(0))\big{)}.$
This follows because for any realization $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $K^{\text{true}}(x)\leq K(\alpha)$ for each $x\in\{0,1\}$ implies that A2 holds, and hence theorem 1 gives $\text{ATE}\in\Theta_{I}(K(\alpha))$. So the inequality follows by monotonicity of probability measures. Next, define $K^{*}(\alpha)\coloneqq\inf\{K\geq 0:p(K,\mathbf{Y}(1),\mathbf{Y}(0))\geq 1-\alpha\}$. This is a non-stochastic, infeasible, “oracle” choice of $K(\alpha)$. We have $K^{*}(\alpha)\leq K(\alpha)$ for all realizations $\mathbf{X}$ (recall that $K(\alpha)$ is random here). That follows because $\underline{p}(K)\leq p(K,\mathbf{Y}(1),\mathbf{Y}(0))$ for all $K$, for all realizations of $\mathbf{X}$ ($\underline{p}$ is also random here) and by definition of $K(\alpha)$. This implies that
$\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K^{*}(\alpha)\text{ for $x\in\{0,1\}$})\leq\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K(\alpha)\text{ for $x\in\{0,1\}$}).$
Finally, by definition, $p(K,\mathbf{Y}(1),\mathbf{Y}(0))\coloneqq\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K\text{ for $x\in\{0,1\}$})$. Thus
$\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K^{*}(\alpha)\text{ for $x\in\{0,1\}$})=p(K^{*}(\alpha),\mathbf{Y}(1),\mathbf{Y}(0))\geq 1-\alpha.$
The last inequality follows by the definition of $K^{*}(\alpha)$, and since $p(\cdot,\mathbf{Y}(1),\mathbf{Y}(0))$ is right continuous. Putting everything together gives $\mathbb{P}_{\text{design}}\big{(}\mathcal{C}\ni\theta(\mathbf{Y}(1),\mathbf{Y}(0))\big{)}\geq 1-\alpha$. This holds for all $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ and therefore the inequality holds for the infimum as well. ∎
Proof of theorem 4.
[…]
  1. 1.
[…]
  2. 2.
[…]
Part 2. Fix $\alpha\in(0,1)$. Since $\Theta_{I}(K)$ is an interval, $\sup_{\theta\in\Theta_{I}(K(\alpha))}|\theta-\text{ATE}_{N}|$ equals
$\max\left\{\left|\big{(}\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big{)}-\text{ATE}_{N}\right|,\;\left|\big{(}\text{LB}_{K(\alpha)}(0)-\text{UB}_{K(\alpha)}(1)\big{)}-\text{ATE}_{N}\right|\right\}.$
[…]
Proof of lemma 1.
By algebra, $\overline{Y}_{z=1}-\overline{Y}_{z=0}=\big{(}\overline{U}_{z=1}-\overline{U}_{z=0}\big{)}+\sum_{i=1}^{N}\beta_{i}\big{(}X_{i}(1)\allowbreak\mathbbm{1}(Z_{i}=1)/N_{1}-X_{i}(0)\mathbbm{1}(Z_{i}=0)/N_{0}\big{)}$ and $\overline{X}_{z=1}-\overline{X}_{z=0}=\big{(}\overline{T}_{1}(a)-\overline{T}_{0}(a)\big{)}+\overline{T}_{1}(c)$. Similarly, we can write $\sum_{i=1}^{N}\beta_{i}\big{(}X_{i}(1)\mathbbm{1}(Z_{i}=1)/N_{1}-X_{i}(0)\mathbbm{1}(Z_{i}=0)/N_{0}\big{)}=\Big{(}\overline{T}_{1}(a)\overline{\beta}_{1}(a)-\overline{T}_{0}(a)\overline{\beta}_{0}(a)\Big{)}+\overline{T}_{1}(c)\overline{\beta}_{1}(c)$. Putting these derivations together gives $\overline{Y}_{z=1}-\overline{Y}_{z=0}=\big{(}\overline{U}_{z=1}-\overline{U}_{z=0}\big{)}+\Big{(}\overline{T}_{1}(a)\overline{\beta}_{1}(a)-\overline{T}_{0}(a)\overline{\beta}_{0}(a)\Big{)}+\overline{T}_{1}(c)\overline{\beta}_{1}(c)$. Dividing by the first stage gives equation (6). ∎
Proof of proposition 2.
Follows immediately from lemma 1. ∎
Proof of theorem 5.
When there are no always takers, equation (6) simplifies to $\text{WaldEstimand}\coloneqq\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}=\frac{\overline{U}_{z=1}-\overline{U}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}+\overline{\beta}_{1}(c)$. Thus we no longer have to worry about balance in always takers, simply because they don’t exist. We only have to worry about balance in $Y_{i}(0)$ across the treatment and control groups. Solving this equation for LATT gives $\overline{\beta}_{1}(c)=\text{WaldEstimand}-(\overline{U}_{z=1}-\overline{U}_{z=0})/\pi$ where recall that $\pi\coloneqq\overline{X}_{z=1}-\overline{X}_{z=0}$ denotes the first stage, and relevance ensures that we are not dividing by zero. This equation and B5 immediately show that the bounds in the theorem statement are valid. Sharpness obtains because the unknown $Y_{i}(0)$ values are completely unconstrained, so long as they satisfy B5. Hence any value in the interval is attainable. ∎

### v3

@@
 We develop a new approach for quantifying uncertainty in finite populations, by using design distributions to calibrate sensitivity parameters in finite population identified sets.
-This yields uncertainty intervals that can be interpreted as identified sets, Bayesian credible sets, or frequentist design-based confidence sets.
-We focus on quantifying uncertainty about the average treatment effect (ATE) due to missing potential outcomes in a randomized experiment, where our approach (1) yields design-based confidence intervals for ATE which allow for heterogeneous treatment effects but do not rely on asymptotics, (2) provides a new motivation for examining covariate balance, and (3) gives a new formal analysis of the role of randomized treatment assignment.
+This yields uncertainty intervals that can be interpreted as identified sets, robust Bayesian credible sets, or uniform frequentist design-based confidence sets.
+We focus on quantifying uncertainty about the average treatment effect, where our approach (1) yields design-based confidence intervals which allow for heterogeneous treatment effects without using asymptotics, (2) provides a new motivation for examining covariate balance, and (3) gives a new formal analysis of the role of randomization.
 We illustrate our approach in three empirical applications.
@@
 1 Introduction
-Since the beginning of statistics, there have been two ways to think about populations:
-  1. 1.
-    They are finite, as in Neyman (1923), or Bowley (1926) who considered “a sample containing $n$ persons or things [that] is selected from a universe containing $N$….” (page 3).
-  2. 2.
-    Or they are infinite, as in Fisher (1930), or Student (1908) who assumed “a random sample has been obtained from an indefinitely large* population….
-*Note that the indefinitely large population need not actually exist” (page 302).
-Inference in finite populations came to be called “design-based inference” or “randomization inference” while inference in infinite populations came to be called “model-based inference” or “super-population inference” (e.g., Wu and Thompson 2020).
-Econometric theory has traditionally focused on super-population inference, going back to Haavelmo (1944) who explained that “when we describe $s$ as a random variable with a certain probability distribution for each fixed set of values of the variables $x$, we are thinking of a class of hypothetical, infinite populations…” (page 51).
-Likewise, much of statistics has also focused on the super-population approach, with survey sampling theory as the main exception (see Little 2004).
-[^3] These distinctions and jargon come from the survey sampling literature (e.g., Särndal et al.
+Many empirical papers use variation across a small number of units to learn about causal effects.
+For example, in many settings the variation is across units like states or industries.
+How does identification of causal effects work in such settings, and how should empirical researchers quantify uncertainty?
+Our paper gives new answers to these questions, focusing on the context of randomized experiments with a finite number of units.
+Traditionally, the $N$ units are assumed to be sampled from a hypothetical infinite “super” population, but recently there has been a resurgence in the “design-based inference” literature (reviewed in section 1.2) which studies explicitly finite populations.
+[^3] The “design-based” versus “super-population” distinction and jargon come from the survey sampling literature (e.g., Särndal et al.
 1978), which is how we use them here.
@@
 We focus on randomized experiments in this paper and hence both meanings of “design-based” apply.
-However, in the past decade there has been a resurgence of interest in design-based inference in explicitly finite populations (see section 1.2).
-Much of this work follows a template largely laid out by Neyman (1923).
-This template has three steps:
-  1. 1.
-    Describe the finite population.
-  2. 2.
-    Describe the design, which explains how the data is obtained.
-  3. 3.
-    Derive properties of the induced design distribution of statistics that depend on the population and the data.
-For example, in the context of survey sampling, the population is a set of numbers $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$ (step 1).
-A classical design is simple random sampling: Sample exactly $n$ units from $\{1,\ldots,N\}$ uniformly at random (step 2).
-Let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be binary random variables indicating which units were sampled.
-A classical theorem is that the sample mean (a function of $\mathbf{S}$ and $\mathbf{Y}$) is design-unbiased for the finite population mean when $\mathbf{S}$ is chosen uniformly at random: $\mathbb{E}_{\text{design}}\left[\frac{1}{n}\sum_{i=1}^{N}S_{i}Y_{i}\right]=\frac{1}{N}\sum_{i=1}^{N}Y_{i}$ (step 3).
-Viewed from the perspective of econometrics, there is a noticeable omission in this template: What about identification?
-Identification analysis is a key step throughout much of super-population based econometrics (e.g., Lewbel 2019 or Molinari 2020); is any of that analysis still useful in finite populations?
-Yes, as we argue in this paper.
-We use ideas from identification theory to construct a new approach for quantifying uncertainty in finite populations, which we call design-based sensitivity analysis.
+With very few exceptions (e.g., Manski and Pepper 2018, Borusyak et al.
+2025, and Rambachan and Roth 2025), most papers that study finite populations do not discuss identification and instead focus solely on proving frequentist properties like consistency or confidence interval validity.
+In contrast, there is a vast and rich literature on identification in infinite populations [^4] See Christensen and Connault (2023), Obradović (2024), Spini (2024), Deaner (2025), Słoczyński (2025), Blandhol et al.
+(2025), and Caetano et al.
+(2026) for a few recent examples, among many others; for comprehensive surveys see Matzkin (2007), Manski (2009), Bontemps and Magnac (2017), Lewbel (2019), Chesher and Rosen (2020), Molinari (2020), and Kline and Tamer (2023).; no corresponding literature on identification in finite populations exists.
+Our first main contribution, therefore, is to develop a formal theory of identification in finite populations.
+The literature on identification in finite populations is so sparse that we are not aware of any source which formally defines identification in an explicitly finite population setting (though some references come close).
+Nonetheless, it is quite common and “conventional” in the prior literature (Heckman and Vytlacil 2007, footnote 6, page 4786) to informally state that a parameter is point identified if there is a known mapping from the distribution of the data induced by randomization (of treatment assignment, for example) into the parameter.
+We state this prior definition formally, and then argue that it has two major problems.
+First, we show that this prior conventional definition relies on hypothetical data that is impossible to observe, since it involves averaging over counterfactual worlds that are incompatible with each other.
+As a consequence, unlike in infinite populations, we prove that this prior conventional definition in finite populations implies the values of every potential outcome for every unit are point identified in any setting where the probability of each treatment value is positive (a type of overlap restriction), including non-experimental settings where treatment assignment depends almost arbitrarily on potential outcomes (this finding builds on ideas in Imbens 2018 and Chen et al.
+2026, as we discuss in section 2.4).
+For example, according to the prior conventional definition, assumptions like unconfoundedness or instrument exogeneity are unnecessary to obtain point identification of treatment effects, so long as an overlap restriction holds.
+Therefore, we argue that the prior conventional definition is not adequate for assessing how informative the actual data one has are about treatment effects.
+Second, the prior conventional definition of identification, as well as the broader literature on finite population design-based inference, pre-supposes that there is some kind of known a priori randomization.
+For example, that randomization may involve treatment assignment, an instrument assignment, or sampling, depending on the setting.
+These may be plausible in randomized experiments or designed surveys, but in many observational settings it is not clear that the variable of interest (like treatment assignment) is randomly chosen, let alone that the distribution it was drawn from is known to the researcher (a point discussed in Manski and Pepper 2018, pages 234–235).
+Consequently, the prior conventional definition of identification in finite populations does not apply to such cases.
+To address these two problems, we develop a different, formal definition of identification for a parameter in a finite population setting.
+Our definition is analogous to the usual definition of identification in infinite population settings (e.g., Manski 2003, Lewbel 2019, Molinari 2020): The finite population identified set for a parameter (such as the average treatment effect, ATE) is the set of parameter values consistent with (a) the finite matrix of data the researcher actually has and (b) any assumptions that they make.
+Our definition directly addresses the two problems with the prior conventional definition described above: First, our definition does not require any knowledge about how treatment is assigned.
+Hence it can be applied in a wider variety of settings.
+Second, it uses actually observed data, rather than hypothetical data that is impossible to observe.
+Relatedly, it does not require any kind of asymptotics or hypothetical infinite populations.
+Consequently, any parameter which cannot be written as a known function of the actually observed data will typically be partially identified.
+The size of the identified set, however, will generally vary with the assumptions and data available, as in the usual infinite population definition of identified sets.
+As a leading example of our identification analysis, we show that the average treatment effect in a randomized experiment is generally partially identified in finite populations.
+[^5] For a recent survey of infinite population approaches to randomized experiments, see Bai et al.
+(2025b).
+This conclusion arises from a subtle distinction between the finite and infinite population settings: Randomization guarantees that the treatment and control groups are exactly balanced in an infinite population, but not in finite populations.
+Since randomization cannot guarantee any level of balance between the treatment and control groups in a finite population, it does not have any identifying power.
+Nonetheless, we show how to use our finite population identified sets combined with assumptions on treatment assignment to obtain a new way to quantify uncertainty in finite populations, which we call design-based sensitivity analysis.
+This is our second main contribution.
+Overall, our method yields uncertainty intervals that can be interpreted as (1) identified sets, (2) robust Bayesian credible sets, or (3) uniform frequentist design-based confidence sets.
 This approach is non-asymptotic, which allows it to be reliably applied in small populations.
-It yields uncertainty intervals that can be interpreted as (1) identified sets, (2) Bayesian credible sets, or (3) frequentist design-based confidence sets.
 And it allows researchers to flexibly impose additional assumptions on unobserved values, transparently quantifying the impact of these assumptions on conclusions.
-We develop this approach in the context of randomized experiments with a binary treatment, where the goal is to quantify the uncertainty in the average treatment effect (ATE) due to missing potential outcomes (we also briefly discuss sampling uncertainty).
-In this context, our approach yields three further contributions: (1) Our uncertainty intervals allow for heterogeneous treatment effects, implying that our approach yields design-based confidence intervals for ATE without relying on asymptotics or imposing homogeneous treatment effects.
-These intervals are not constructed via hypothesis test inversion, but rather are based on ideas from partial identification.
-(2) We give a new motivation for examining covariate balance.
+We summarize our approach in section 1.1.
+In particular, focusing on their frequentist interpretation, our uncertainty intervals avoid several limitations of the confidence intervals that are currently available in the literature: (i) They do not rely on large-$N$ asymptotics, which can perform poorly if $N$ is small (e.g., $N=10$) and (ii) they allow for heterogeneous treatment effects.
+They are also valid uniformly over the set of all possible matrices of potential outcomes consistent with the assumptions (an important feature in partially identified settings; see Canay and Shaikh 2017).
+Furthermore, unlike the confidence intervals in the literature, our uncertainty intervals have both Bayesian and identification-based interpretations.
+This allows empirical researchers to choose their most preferred interpretation.
+This interpretational flexibility is one benefit obtained from a quantification of uncertainty that begins with identification.
+In particular, our uncertainty intervals are not based on hypothesis test-inversion, but are instead based on ideas from partial identification.
+As mentioned above, we develop this approach in the context of randomized experiments with a binary treatment, where the goal is to quantify the uncertainty in the average treatment effect due to missing potential outcomes.
+We extend the analysis in three main directions: In section 5 we give a new motivation for examining covariate balance.
 Specifically, by making an assumption that explicitly links covariates to potential outcomes, we show that observed imbalances in covariates have identifying power for the unobserved realized imbalance in potential outcomes, which allows researchers to obtain tighter bounds on ATE.
-And (3) we give a new formal analysis of the role of randomized treatment assignment.
+In section 6 we study instrumental variable based analysis under one-sided noncompliance.
+And in section 7.1 we discuss uncertainty due to non-sampled units.
 1.1 Our Approach to Quantifying Uncertainty in Experiments: A Practitioner’s Guide
 Here we briefly sketch our approach, with an eye toward practical take-aways.
+We focus on causal inference here; see appendix LABEL:sec:sampleSurveyGuide for a parallel exposition in the context of survey sampling.
 Consider the population of six units in table 1 (left).
@@
 Figure 1: Example output for a design-based sensitivity analysis, $N=10$. See section 8.1 for data details.
-Finally, in step 3, we recommend combining these two plots as follows: For any desired ex ante probability $1-\alpha\in(0,1)$, use the bottom plot to find the smallest magnitude of imbalance that occurs with at least $100(1-\alpha)$ex ante probability; denote this value by $K(\alpha)$.
-For example, if we set $0.91-\alpha=$ 0.9 then the figure shows us that $K=18.9$ satisfies $\underline{p}(K)=1-\alpha$ (see the dotted lines).
+Finally, in step 3, we recommend combining these two plots as follows: For any desired ex ante probability $1-\alpha\in(0,1)$, use the bottom plot to find the smallest magnitude of imbalance that occurs with at least $100(1-\alpha)$% ex ante probability; denote this value by $K(\alpha)$.
+For example, if we set $1-\alpha=0.9$ then the figure shows us that $K=18.9$ satisfies $\underline{p}(K)=1-\alpha$ (see the dotted lines).
 We can then move to the top plot to read off bounds on ATE for this value of the sensitivity parameter, as shown in the dotted lines.
-In this case the bounds are about $[-6,22]$.
+In this case the bounds are about [6,22][-6,22].
 Figure 1(b) shows these bounds for all values of $1-\alpha$.
+[^6] Note: For $1-\alpha$ close to zero but strictly positive, the bounds in figure 1(b) are non-singleton.
+This arises since exact balance is impossible in the worst case completions of table 1; this is why the plot of $\underline{p}$ in figure 1(a) is flat at zero up to a strictly positive value of $K$.
+We discuss this point further in section 2.3.
 We show that these uncertainty intervals have three interpretations:
@@
     Third, for any fixed $1-\alpha$, we show that these bounds are design-based confidence intervals.
-That is, across repeated re-assignments of treatment, these bounds will contain the true ATE at least $100(1-\alpha)$of the time.
+That is, across repeated re-assignments of treatment, these bounds will contain the true ATE at least $100(1-\alpha)$% of the time.
 Overall, we recommend that researchers report plots like figure 1(b), which can then be interpreted any of these ways, depending on personal preference.
@@
 Manski and Pepper (2018) derive identified sets in an explicitly finite population setting.
-Like them, we also consider the identifying power of bounded variation type assumptions, which restrict just how far potential outcomes can be from each other.
-The main difference is that they focus on a setting with observational data whereas we consider randomized experiments.
-Our focus on experiments allows us to use randomization itself to calibrate the magnitude of the bounded-variation sensitivity parameter, which is the design-based sensitivity analysis we develop in this paper.
+Like them, we also consider the identifying power of bounded variation type assumptions.
+A main difference is that they focus on a setting with observational data whereas we primarily consider randomized experiments.
+Our focus on experiments allows us to use randomization itself to calibrate the magnitude of the bounded-variation sensitivity parameter $K$.
 Like us, Greenland and Robins (1986) also describe the problem of drawing conclusions in finite populations as an identification problem which must be solved by making assumptions about unobserved values of variables.
@@
 That paper as well as the rest of this prior literature, however, is largely verbal discussion.
-One of our main contributions is to use those informal ideas to provide a new, detailed, and formal approach to quantifying uncertainty in small finite populations.
+One of our main contributions is to build on those observations to provide a new, detailed, and formal approach to identification and quantifying uncertainty in finite populations.
 The standard approach to uncertainty quantification in finite populations is to construct frequentist design-based confidences sets.
@@
 (2020, 2023), Hong et al.
-(2020), Rambachan and Roth (2020), Wu and Ding (2021), Eckles et al.
+(2020), Wu and Ding (2021), Eckles et al.
 (2020), Imbens and Menzel (2021), Zhao and Ding (2021), Sävje (2021), Bojinov et al.
-(2021), Xu (2021), Xu and Wooldridge (2022), Athey and Imbens (2022), Roth and Sant’Anna (2023), Pollmann (2023), Wooldridge (2023), Startz and Steigerwald (2023), Borusyak et al.
-(2024), and Sancibrián (2024), among many others.
+(2021), Xu (2021), Xu and Wooldridge (2022), Athey and Imbens (2022), Roth and Sant’Anna (2023), Pollmann (2023), Wooldridge (2023), Startz and Steigerwald (2023), Sancibrián (2024), Borusyak et al.
+(2025), Rambachan and Roth (2025), and Bai et al.
+(2025a), among many others.
 Also see Imbens and Rubin (2015) and Ding (2024) for book level surveys.
-This literature largely follows the template of Neyman (1923) that was discussed earlier, and does not do explicit identification analysis.
+This literature largely follows a template laid out by Neyman (1923) that jumps straight to deriving various frequentist properties, without doing any identification analysis.
+Two exceptions are Borusyak et al.
+(2025) and Rambachan and Roth (2025) which both discuss identification, based on the conventional definition, and do asymptotics-based inference (section 4.3 of Borusyak et al.
+2025 also discuss exact inference under a constant effects assumption).
+Our finding that all potential outcomes are point identified under the conventional definition builds on ideas in Imbens (2018) and Chen et al.
+(2026), as we discuss in section 2.4.
+Rather than further studying identification, Chen et al.
+(2026) derive various frequentist testing and Bayesian updating implications of the conventional definition of identification, concluding that this conventional definition “translates to little, if any, practical learning” (page 14).
 2 Finite Population Identification
-We first define the finite population identified set. We then use this definition to derive identified sets for finite population average treatment effects under various assumptions.
+We begin by describing our approach to identification: We define the finite population identified set in section 2.1.
+We then use this definition to derive identified sets for finite population average treatment effects under various assumptions in section 2.2, and we discuss the role of randomization for identification in section 2.3.
+We explain our concerns with the prior definition of identification in section 2.4.
 2.1 The Finite Population Identified Set
-Consider the standard potential outcomes model with a binary treatment. There are four components required to define an identified set:
+Consider the standard potential outcomes model with a binary treatment. There are four components required to conduct an identification analysis:
 1.
 The population: Suppose there are $N$ units in our population.
-Let =$\mathcal{I}=\{1,\ldots,N\}$ be the set of indices for these units.
+Let $\mathcal{I}=\{1,\ldots,N\}$ be the set of indices for these units.
 Each unit $i\in\mathcal{I}$ is associated with the vector of numbers $(Y_{i}(1),Y_{i}(0),X_{i},W_{i})$, where $Y_{i}(1)$ and $Y_{i}(0)$ are potential outcomes, $X_{i}\in\{0,1\}$ is a realized binary treatment, and $W_{i}$ is a $d_{W}$-vector of covariates.
 The population is the $N\times(3+d_{W})$ matrix of all of these numbers.
-Let $\mathbf{P}^{\text{true}}\coloneqq(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{X},\mathbf{W})$ denote this matrix.
+Let $\mathbf{P}^{\text{true}}\coloneqq(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{X},\mathbf{W})$ denote this population matrix.
 We use $\mathbf{P}$ to denote alternative possible values of this population matrix.
@@
 2.
-Assumptions: We do not observe $\mathbf{P}^{\text{true}}$.
-We will make assumptions about it, however.
+Assumptions: We do not observe all elements of $\mathbf{P}^{\text{true}}$ because of the fundamental problem of causal inference.
+We typically make assumptions about its missing elements, the missing potential outcomes.
 Formalize these assumptions as the restriction that $\mathbf{P}^{\text{true}}\in\mathcal{P}$ where $\mathcal{P}$ is a known set of $N\times(3+d_{W})$ matrices.
@@
 The data: Finally, we must describe the data that is observed to the econometrician.
-To allow for sampling, let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be a vector of indicators where $S_{i}=1$ if unit $i$ appears in the dataset, and zero otherwise, as in section 1.
+To allow for sampling, let $\mathbf{S}=(S_{1},\ldots,S_{N})$ be a vector of indicators where $S_{i}=1$ if unit $i$ appears in the dataset, and zero otherwise.
 Let $\mathbf{Y}=(Y_{1},\ldots,Y_{N})$ where $Y_{i}\coloneqq Y_{i}(1)X_{i}+Y_{i}(0)(1-X_{i})$ for all $i\in\mathcal{I}$.
@@
 $\Theta_{I}\coloneqq\{\theta\in\Theta:\theta=\theta(\mathbf{P})\text{ for some }\mathbf{P}\in\mathcal{P}\text{ such that }\text{MakeData}(\mathbf{P},\mathbf{S})=\mathbf{P}^{\text{data}}\}.$
-This definition follows the usual description of the identified set as “the set of parameters…that are consistent with the model and the data” (Tamer 2010, page 184), but with two variations to clarify what is meant by “data” when the population is finite: First, even when all units are sampled, the data is defined by a finite dimensional matrix rather than a joint distribution of random variables (as in def.
+This definition follows the usual description of the identified set as “the set of parameters…that are consistent with the model and the data” (Tamer 2010, page 184), but with close attention to what is meant by “data”: In a finite population, the “data” is defined by a finite dimensional matrix rather than a joint distribution of random variables (as in def.
 3.1 on page 5324 of Matzkin 2007, for example; also see our appendix LABEL:sec:popDistributions for a related discussion).
-Second, our definition allows for data on some units to be missing (i.e., not sampled), by setting $S_{i}=0$.
-This is perhaps unusual, since uncertainty due to sampling is traditionally separated from other types of uncertainty, as in Koopmans’ (1949, page 132) original definition of identification.
-Nonetheless, as we show in this paper, definition 1 provides a useful foundation to quantify uncertainty from both missing units and missing potential outcomes.
-In particular, $\Theta_{I}$ has two key features: (1) It can always be computed with the data the analyst actually has at hand, and (2) like traditional definitions of identification, it is guaranteed to contain the true parameter of interest, so long as the model is not false.
+Our definition also allows for data on some units to be missing (i.e., not sampled, so that $S_{i}=0$); we discuss this further in section 7.1.
+$\Theta_{I}$ has two key features: (1) It can always be computed with the data the analyst actually has at hand, and (2) like the usual definition of the identified set in an infinite population, it is guaranteed to contain the true parameter of interest, so long as the model is not false.
 We will revisit these properties below.
-Finally, note that some authors define the “data” as the design distribution of the observed data, rather than the dataset that is actually observed.
-This leads to a different definition of identification that can be applied to finite populations; for example, see the definition of “sampling identification” in Florens and Simoni (2021, def 2.1 on page 5).
-This design distribution typically depends on the entire population matrix and hence is never actually knowable, even if all units are sampled (because not all potential outcomes are observed).
-Moreover, under this alternative definition of identification, the existence of a design-unbiased estimator for a parameter implies that the parameter is point identified.
-In contrast, under definition 1, any parameter that depends on at least one unknown value, like a missing potential outcomes, will typically be partially identified, even if there is a design-unbiased estimator for it.
-In this paper we argue that this feature of our definition leads to a useful approach for quantifying uncertainty in finite populations.
 2.2 Identified Sets for ATE
-To simplify the exposition, from here on we assume that all units in the population are observed, and focus solely on the uncertainty due to missing potential outcomes.
-In section 7.1 we show that the extension to incorporate sampling is straightforward.
+To simplify the exposition, we generally assume that all units in the population are observed ($S_{i}=1$ for all $i\in\mathcal{I}$).
+We show that the extension to incorporate sampling is straightforward in section 7.1.
 We also focus on the identification of the average treatment effect for brevity.
 We briefly discuss other parameters in section 7.3.
-It is well known that, because averages are influenced by the values of outlier observations, bounds on ATE are usually infinite without some kind of restriction on the magnitude of outliers.
-Hence we maintain the following assumption throughout the paper.
+It is well known that bounds on mean based parameters are usually infinite without some kind of restriction on the values they can take.
+Hence we maintain the following assumption throughout the paper, which is standard in the literature on partial identification.
 Assumption A1 (Bounded outcomes).
-There are known values $-\infty<y_{\text{min}}<y_{\text{max}}<\infty$ such that $Y_{i}(x)\in[y_{\text{min}},y_{\text{max}}]$ for all $i\in\mathcal{I}$, for each $x\in\{0,1\}$.
+There are known values $\infty<y_{\text{min}}<y_{\text{max}}<\infty$ such that $Y_{i}(x)\in[y_{\text{min}},y_{\text{max}}]$ for all $i\in\mathcal{I}$, for each $x\in\{0,1\}$.
 In some applications the values of $y_{\text{min}}$ and $y_{\text{max}}$ can be set to their logical values, such as 0 to 100 for test scores.
@@
 We discuss several variations on this assumption in section 7.2, including an assumption that does not require selecting a priori bounds on outcomes.
-A1 alone has identifying power for ATE, via a finite population version of the Manski (1990) bounds.
-This result is a special case of a more general result that uses the following exogeneity assumption.
+A1 alone has identifying power for ATE, via a finite population version of the Manski (1990) bounds. Those bounds are a special case of theorem 1 below, which uses the following assumption.
 Assumption A2 ($K$-approximate mean balance).
@@
 Theorem 1.
-Suppose A1 and A2 hold, and $\mathbf{P}^{\text{data}}$ is known. Then, for each $x\in\{0,1\}$, the identified set for $\overline{\mathbf{Y}(x)}$ is $[\text{LB}_{K}(x),\text{UB}_{K}(x)]$ where
+Suppose A1 and A2 hold, and $\mathbf{P}^{\text{data}}$ is known. Then, for each $x\in\{0,1\}$, the identified set for $\overline{\mathbf{Y}(x)}$ is $\text{LB}_{K}(x),\text{UB}_{K}(x)]$ where
 $\text{LB}_{K}(x)$ $\coloneqq\overline{Y}_{x}\frac{N_{x}}{N}+\max\{y_{\text{min}},\overline{Y}_{x}-K\}\frac{N_{1-x}}{N},\text{ and}$
@@
 2.3 Randomization and Identification in Finite Populations
-[…]
-$\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)=\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)$
-for $x\in\{0,1\}$.
-In fact, for many values of the vectors $\mathbf{Y}(x)=(Y_{1}(x),\ldots,Y_{N}(x))$, it is impossible for exact balance to hold regardless of the values of realized treatment $\mathbf{X}=(X_{1},\ldots,X_{N})$, a point noted by Greenland and Robins (1986, page 415).
-So statistical independence is not an appropriate formalization of random assignment in finite populations.
-Instead, we make the following assumption.
+In practice, researchers often motivate assumptions that two groups are in some sense comparable by appealing to random assignment.
+So consider the following common formalization of random assignment in finite populations.
 Assumption A3 (Random assignment).
-The size of the treatment and control groups, $N_{1}$ and $N_{0}$, are fixed a priori.
+The size of the treatment and control groups, $N_{1}$ and $N_{0}$, are fixed a priori, with $0<N_{1}<N$.
 $\mathbf{X}=(X_{1},\ldots,X_{N})$ is a single realization from the known probability distribution $\mathbb{P}_{\text{design}}$ on $\{0,1\}^{N}$ defined by
@@
 As has long been recognized, randomization in finite populations does not guarantee any level of ex post balance (e.g., Greenland 1990).
-The identified set, by definition, must contain the true parameter whenever the model is true.
+The identified set, by definition, must contain the true parameter value whenever the model assumptions are true.
 Consequently, for any finite population, randomization has no identifying power.
@@
 Put differently, after treatment has been assigned, we cannot rule out the possibility that potential outcomes are substantially imbalanced across the treatment and control groups, even if it was unlikely a priori.
-Thus, from an identification perspective, randomization does not shrink the domain-of-consensus bounds in finite populations.
-Nonetheless, in the next section we’ll show that theorem 2 provides an overly pessimistic view of the value of randomization.
-There we reinterpret randomization as a procedure that affects our beliefs about ex post balance.
+Thus the assumption that treatment was randomly assigned does not rule out any values of the unknown potential outcomes.
+Hence it does not shrink the domain-of-consensus bounds in finite populations.
+Nonetheless, in section 3 we will reinterpret randomization as a procedure that affects our beliefs about ex post balance.
 Specifically, we will use randomization to assess the plausibility of a specific choice of the sensitivity parameter $K$.
 This will let us use the identification result in theorem 1 to perform a sensitivity analysis motivated by random assignment of treatment.
+That analysis will yield uncertainty intervals that have interpretations as identified sets, robust Bayesian credible sets, or uniform frequentist confidence sets.
+Exact Balance in Infinite versus Finite Populations
+[…]
+This qualitative distinction between the finite and infinite setting explains why theorem 2 comes to a different conclusion about identification of ATE from traditional infinite population identification analyses: In the infinite population setting, exact balance is assumed to hold and hence ATE is point identified.
+But in the finite population setting, exact balance is not guaranteed even under random assignment, and is sometimes even logically impossible, and hence ATE is partially identified.
+That said, because any level of approximate balance is more likely when $N$ is large, there is a connection between these two results, which we study in section 3.3.
+Also see section 7.3, where we discuss various measures of balance.
+2.4 The Prior Definition of Finite Population Identification
+Now that we have explained our approach to identification, we turn to examine the prior definition.
+First, what is the prior definition?
+We are not aware of any books or papers that explicitly and formally define identification for finite populations.
+For example, Borusyak et al.
+(2025) discuss identification but do not provide a formal definition, and as representative of the rest of the literature, their first formal result is a consistency proof, not an identification theorem.
+The book by Tillé (2020) is an extensive and detailed survey of sampling from finite populations, but does not discuss identification.
+The textbook by Ding (2024) discusses causal inference in finite populations in detail, starting from chapter 2, but a definition of identification does not appear until chapter 10, only after shifting focus from randomized experiments to observational data.
+That definition (10.1 on page 128) states that “a parameter $\theta$ is nonparametrically identifiable if it can be written as a function of the distribution of the observed data without any parametric model assumptions”.
+Similar definitions appear in other references, going back to the earliest sources (such as Koopmans 1949, Koopmans and Reiersøl 1950, Hurwicz 1950, Rothenberg 1971).
+Verbal definitions like these are ambiguous about the meaning of “the distribution of the observed data” (Ding 2024) or the “joint distribution of the observations” (Koopmans 1949, page 125).
+There are different ways to formalize such statements and they lead to substantially different definitions of identification.
+As we argue below, this issue is particularly important in the finite population setting.
+Our definition 1 is one such formalization, where by “data” we mean the matrix $\mathbf{P}^{\text{data}}$ that the researcher actually has at hand.
+In this case the “distribution” of this data can be formalized as a specific discrete distribution (see appendix LABEL:sec:popDistributions for details).
+However, based on the informal discussions in the literature, and our personal communication with researchers in this field, “data” is conventionally interpreted as the design distribution of the observed data, rather than the dataset that is actually observed.
+This interpretation leads to the following definition of point identification for finite populations.
+Definition 2 (Conventional).
+Let $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ be the $N\times(2+d_{W})$ matrix of potential outcomes and covariate values for all units in the population $\mathcal{I}=\{1,\ldots,N\}$.
+Let $\theta(\cdot)$ be a functional defined on the set of logically possible values of this matrix.
+Let $\theta^{\text{true}}\coloneqq\theta((\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W}))$ denote its true value.
+For each $i\in\mathcal{I}$, let $X_{i}^{\text{new}}\in\{0,1\}$ and $S_{i}^{\text{new}}\in\{0,1\}$ be random variables.
+Let $\mathbf{X}^{\text{new}}\coloneqq(X_{1}^{\text{new}},\ldots,X_{N}^{\text{new}})$ and $\mathbf{S}^{\text{new}}\coloneqq(S_{1}^{\text{new}},\ldots,S_{N}^{\text{new}})$.
+Let $\mathbb{P}_{\text{design}}$ denote the joint distribution of $(\mathbf{X}^{\text{new}},\mathbf{S}^{\text{new}})$; it is called the design distribution.
+Let $Y_{i}^{\text{new}}\coloneqq Y_{i}(X_{i}^{\text{new}})$.
+Let $\mathcal{S}=\{i\in\mathcal{I}:S_{i}^{\text{new}}=1\}$ denote the random set of sampled indices.
+Let $\mathcal{D}\coloneqq\{(Y_{i}^{\text{new}},X_{i}^{\text{new}},W_{i})\in\mathbb{R}^{2+d_{W}}:i\in\mathcal{S}\}$.
+Say $\theta^{\text{true}}$ is point identified if there is a known mapping from the probability distribution of the random set $\mathcal{D}$ to $\theta^{\text{true}}$.
+We believe definition 2 accurately formalizes the conventional definition of finite population identification in the prior literature.
+For example, Heckman and Vytlacil (2007, footnote 6, page 4786) say “identification in small samples requires establishing the sampling distribution of estimators, and adopting bias as the criterion for identifiability” (see our proposition 1 below) and that “this approach is conventional in classical statistics,” although they do not give a formal definition which we could compare against.
+This definition also resembles the definition of “sampling identification” in Florens and Simoni (2021, definition 2.1 on page 5), although that paper is not about design-based inference or finite populations.
+Definition 2 has the following seemingly innocuous and familiar implication.
+Proposition 1.
+Consider the setting of definition 2.
+Suppose there is a known function $m$ such that $\theta^{\text{true}}$ is the unique solution to $\mathbb{E}_{\text{design}}[m(\mathcal{D},\theta)]=0$.
+Then $\theta^{\text{true}}$ is point identified.
+This result says that a parameter is point identified if it is the unique solution to a set of design-based moment conditions.
+A special case arises when there is an $m$ is such that we can write $\theta^{\text{true}}=\mathbb{E}_{\text{design}}[m(\mathcal{D})]$.
+In this case the function $m(\mathcal{D})$ is usually called a design-unbiased estimator of $\theta^{\text{true}}$.
+Moment-based reasoning like in proposition 1 is used in the identification analysis of Borusyak et al.
+(2025, page 88, first paragraph of section 3.1) and Rambachan and Roth (2025, page 13, first displayed equation in second column).
+Proposition 1, however, implies the following result.
+Proposition 2.
+Consider the setting of definition 2. Suppose $\mathbb{P}_{\text{design}}(S_{i}^{\text{new}}=1)=1$ for all $i\in\mathcal{I}$. Then, according to definition 2, the following hold.
+  1. 1.
+    $p_{i}\coloneqq\mathbb{P}_{\text{design}}(X_{i}^{\text{new}}=1)$ is point identified for all $i\in\mathcal{I}$.
+  2. 2.
+    $Y_{i}(1)$ is point identified for each $i\in\mathcal{I}$ with $p_{i}>0$. $Y_{i}(0)$ is point identified for each $i\in\mathcal{I}$ with $p_{i}<1$.
+Proof of proposition 2.
+Part 1 follows from $p_{i}=\mathbb{E}_{\text{design}}(X_{i}^{\text{new}})$ and proposition 1. For part 2, define
+$\widehat{Y_{i}(1)}\coloneqq\frac{X_{i}^{\text{new}}Y_{i}^{\text{new}}}{p_{i}}\qquad\text{and}\qquad\widehat{Y_{i}(0)}\coloneqq\frac{(1-X_{i}^{\text{new}})Y_{i}^{\text{new}}}{1-p_{i}}.$
+Then
+$\mathbb{E}_{\text{design}}[\widehat{Y_{i}(1)}]=\frac{1}{p_{i}}\mathbb{E}_{\text{design}}[X_{i}^{\text{new}}Y_{i}(X_{i}^{\text{new}})]\\ =\frac{1}{p_{i}}\mathbb{E}_{\text{design}}[X_{i}^{\text{new}}]Y_{i}(1)\\ =Y_{i}(1).$
+Similarly, $\mathbb{E}_{\text{design}}[\widehat{Y_{i}(0)}]=Y_{i}(0)$. Apply proposition 1. ∎
+We assumed all units are sampled in proposition 2 for simplicity, but the result can be generalized to allow for sampling.
+Proposition 2 combines ideas from Imbens (2018, footnote 3) and Chen et al.
+(2026).
+Specifically, using a different proof strategy and restricting attention to the binary outcome case, Chen et al.
+(2026) showed that the proportion of units who benefit from treatment is point identified, according to the conventional definition 2 of identification in finite populations.
+Imbens (2018) considered a randomized experiment with $N=1$ and $\mathbb{P}(X_{1}^{\text{new}}=1)=0.5$ and showed that there exists a design-unbiased estimator of that units’ unit-level treatment effect, $Y_{1}(1)-Y_{1}(0)$.
+Our proposition 2 builds on these ideas to show that all potential outcomes are point identified, according to definition 2, as long as the probability of treatment is strictly between zero and one for all units.
+Proposition 2 applies to randomized experiments, which typically guarantee $p_{i}\in(0,1)$ by design (e.g., A3 holds), but it also applies to many observational settings as well, because this result does not assume that $p_{i}$ is known for all $i\in\mathcal{I}$.
+For example, unconfoundedness restricts the treatment assignment probabilities $p_{i}$ to be functions of observed covariates (in which case it is called the propensity score).
+Proposition 2 implies that such a priori restrictions on treatment probabilities are unnecessary to point identify potential outcomes, according to definition 2; it could even be that the unit-level treatment assignment probabilities $p_{i}=f(Y_{i}(1),Y_{i}(0))$ are functions of potential outcomes themselves; all that matters is that the ex ante probability of treatment is not degenerate on 0 or 1.
+Put differently, unconfoundedness is unnecessary according to the conventional definition of identification; overlap ($p_{i}\in(0,1)$ for all $i\in\mathcal{I}$) is sufficient for point identification of all unit-level treatment effects.
+Similarly, the conventional definition implies that all unit-level treatment effects are point identified for compliers in an instrumental variables setting, even if the instrument is endogenous (see appendix LABEL:sec:conventionalAppendix for details).
+Essentially all of these conclusions implied by proposition 2 differ from the conclusions of classical infinite population identification theory.
+The fact that a parameter can be partially identified under infinite population theory but point identified according the conventional definition was first pointed out by Chen et al.
+(2026), who focused on the proportion of units who benefit from treatment with a binary outcome.
+[^8] They use this finding to motivate a careful study of the implications of the conventional definition of identification for Bayesian updating and various frequentist concepts, rather than exploring different definitions of identification as we do.
+In a randomized experiment, that parameter is generally partially identified in infinite population theory (Makarov 1982, Manski 1997, Fan and Park 2010), and yet the conventional definition 2 implies that it is point identified.
+As they note, in the binary treatment and binary instrument setting of Imbens and Angrist (1994), this implies that the proportion of compliers is point identified even without a no-defiers assumption, another inconsistency with infinite population theory.
+Our result goes further to show that the compliance type (i.e., complier, defier, always taker, or never taker) is point identified for every unit, according to the conventional definition of identification (see appendix LABEL:sec:conventionalAppendix), a further inconsistency with infinite population theory.
+Discussion: Averaging over Units or over Counterfactual Worlds
+The different conclusions implied by different definitions of identification are a direct consequence of how “data” is defined.
+The conventional definition 2 uses data that is not even in principle available to real-world researchers.
+That follows since means like $\mathbb{E}_{\text{design}}$ average over hypothetical, counterfactual worlds that are mutually incompatible with each other—such as a world where unit $i$ is treated and a world where unit $i$ is not treated.
+By allowing identification to depend on information from both such worlds, this definition allows one to conclude that both $Y_{i}(1)$ and $Y_{i}(0)$ are point identified.
+This conclusion suggests that there is, in fact, no fundamental problem of causal inference (Holland 1986).
+In stark contrast, both our definition 1 and the standard definition of infinite population identification do not assume one can observe “data” from such mutually incompatible worlds.
+This holds for our definition 1 since the identified set only depends on $\mathbf{P}^{\text{data}}$, the matrix of data that is actually observed by real empirical researchers.
+The standard infinite population definition is also not based on data from incompatible worlds, and hence does not lead to results like proposition 2.
+Specifically, consider a joint distribution of random variables $(Y,X)$, representing the population “data” of realized outcomes and treatments.
+As explained in Heckman and Vytlacil (2007, section 2.2), for example, each realization $(Y(\omega),X(\omega))$ from this distribution represents a single unit $\omega$ in some set of units $\Omega$ (they assume $\Omega=[0,1]$) who is either treated ($X(\omega)=1)$ or not ($X(\omega)=0$).
+Consequently, even though the population is infinite, each unit is only treated once.
+The “data” available to researchers comes from a single assignment of each of the infinitely many units to be treated or not.
+Concretely, infinite population expectations like $\mathbb{E}\left[\frac{YX}{\mathbb{P}(X=1\mid W)}\right]$, as used in a selection-on-observables analysis (e.g., Imbens and Rubin 2015, section 12.4), do not involve averaging worlds where some units are treated with other worlds where those same units are not treated.
+Rather, such averages are over different units in the infinite population instead of over different counterfactual worlds involving different treatment assignments, as in $\mathbb{E}_{\text{design}}$.
+In summary, these issues with the conventional definition motivate our development of a different theory of identification, based on definition 1.
 3 Design-Based Sensitivity Analysis
+Thus far we have defined and studied identification in finite populations.
+We also derived the identified set for ATE, $\Theta_{I}(K)$, under the $K$-approximate mean balance assumption.
+In this section we show how to use this identified set, combined with a known design distribution of treatment, to quantify uncertainty about ATE.
 3.1 A Design-Based Approach to Calibrating $K$
@@
 Proposition 3.
-Suppose A1 holds.
+Suppose A1 and A3 hold.
 Then for any realization $\mathbf{X}$, $\underline{p}$ is monotonic, $\underline{p}$ is right continuous, $\lim_{K\rightarrow 0}\;\underline{p}(K)=0$, and $\lim_{K\rightarrow\infty}\;\underline{p}(K)=1$.
 $K(\alpha)$ is therefore the smallest value of $K$ such that $K^{\text{true,max}}\leq K$ occurs with probability at least $1-\alpha$, according to the “robust” prior $\underline{p}$, which implies the same would be true for any other prior consistent with our baseline assumptions about the potential outcomes.
-So $[0,K(\alpha)]$ is a Bayesian credible set for $K^{\text{true,max}}$ with respect to $\underline{p}$; it is a set which contains $K^{\text{true,max}}$ with probability about $1-\alpha$ (the true probability is slightly higher than $1-\alpha$ since $\underline{p}$ is not continuous).
+So $0,K(\alpha)]$ is a Bayesian credible set for $K^{\text{true,max}}$ with respect to $\underline{p}$; it is a set which contains $K^{\text{true,max}}$ with probability about $1-\alpha$ (the true probability is slightly higher than $1-\alpha$ since $\underline{p}$ is not continuous).
 And, again, this would also be true for any other prior consistent with our baseline assumptions.
@@
 Keep in mind that $\mathbf{X}^{\text{new}}$ is the only random vector here.
-Theorem 3 shows that $\Theta_{I}(K(\alpha))$ is a $100(1-\alpha)$design-based confidence set.
+Theorem 3 shows that $\Theta_{I}(K(\alpha))$ is a $100(1-\alpha)$% design-based confidence set.
 That is, across repeated random assignments of treatment, it will contain the true parameter with probability at least $1-\alpha$.
@@
   1. 1.
-[…]
+    If $\mu(1)-\mu(0)\neq 0$, $\underline{p}(K^{\text{bp}})\xrightarrow{p}1$ as $N\rightarrow\infty$.
   2. 2.
     Let $\text{ATE}_{N}$ denote the finite population ATE. For any $\alpha\in(0,1)$,
-[…]
+    $\sup_{\theta\in\Theta_{I}(K(\alpha))}|\theta-\text{ATE}_{N}|\xrightarrow{p}0\qquad\text{as $N\rightarrow\infty$.}$
 Theorem 4 shows that a design-based sensitivity analysis will eventually yield very precise conclusions in a large population, whereas an analysis based solely on identification does not.
@@
 $p$-values are commonly interpreted as quantitative measures of the evidence against the null hypothesis, with small values interpreted as “strong evidence” against the null and larger values interpreted as weaker evidence against it.
-This is a controversial interpretation (e.g., see the entire 2019 special issue of The American Statistician on “a world beyond $p<0.05$”).
+This is a controversial interpretation (e.g., see the entire 2019 special issue of The American Statistician on “a world beyond p<0.05p<0.05”).
 In light of this debate about $p$-values, our measure $\underline{p}(K^{\text{bp}})$ can be viewed as an alternative summary statistic that has a well justified interpretation as a quantitative measure of how much evidence the data provides for a specific conclusion.
@@
 We generated a nested sequence of five populations, with $N\in\{10,20,40,100,400\}$.
-These populations have $[y_{\text{min}},y_{\text{max}}]=[0,1]$ and heterogeneous treatment effects such that $\text{ATE}_{N}$ converges to 0.25 as $N$ gets large.
+These populations have $y_{\text{min}},y_{\text{max}}]=[0,1]$ and heterogeneous treatment effects such that $\text{ATE}_{N}$ converges to 0.25 as $N$ gets large.
 We assigned treatment via A3 such that $N_{1}=N/2$ for all $N$ while ensuring the populations are nested.
@@
 Appendix LABEL:sec:numericalAppendix gives further details on how we generated the data.
-[…]
+First we demonstrate the convergence results of theorem 4.
+Its first part shows that $\underline{p}(K^{\text{bp}})\xrightarrow{p}1$ as $N\rightarrow\infty$.
+In the proof, we showed that $\underline{p}(\cdot)$ converges to a step function at zero.
+Figure 2(a) demonstrates this convergence, which implies convergence of $\underline{p}(K^{\text{bp}})$ to 1 so long as the limiting breakdown point is nonzero.
+The second part of theorem 4 shows that, for any $\alpha\in(0,1)$, the distance between the set $\Theta_{I}(K(\alpha))$ and the ATE value obtained under point identification ($K=0$) converges to zero as $N\rightarrow\infty$.
+Figure 2(b) plots $\Theta_{I}(K(\alpha))$ as a function of $1-\alpha$, for the five different values of $N$.
+The bounds are centered at the naive difference-in-means estimand $\overline{Y}_{1}-\overline{Y}_{0}$, which varies with $N$ but converges to 0.25 by construction.
+Again, consistent with the theory, we see that for any $\alpha\in(0,1)$ our bounds shrink as $N$ gets larger.
+This plot also shows the first part of theorem 4, since $\underline{p}(K^{\text{bp}})$ is the point at which the bounds intersect the horizontal axis at zero.
+We see that this point converges to 1 as $N$ gets large.
+Indeed, for $N=100$ our bounds are strictly positive for almost all values of $1-\alpha$.
 (a) Convergence of $\underline{p}$ to a step function at zero as population size $N$ increases.
@@
 subject to (i) the data constraints that $Y_{i}(X_{i})=Y_{i}$ for all $i=1,\ldots,N$, (ii) the approximate $K$-means balance constraint
-$-K\leq\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)\leq K,$
+$K\leq\frac{1}{N_{1}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=0)\leq K,$
 and (iii) the predictive covariates constraint $R_{Y(x)\sim q(W)}^{2}\geq\lambda$, which is equivalent to
@@
 We then study finite population identification under $K$-approximate mean balance type assumptions.
-In the special case of one-sided noncompliance, we show that the finite population identified set for the realized LATT has a simple form that is analogous to classical large population results.
+In the special case of one-sided noncompliance, we show that the finite population identified set for the realized LATT has a simple form that is analogous to classical infinite population results.
 We then show how to use this result to do design-based sensitivity analysis.
-In all of this analysis we allow for heterogeneous treatment effects (in contrast to some prior work on design-based instrumental variable analysis, such as Rosenbaum 1996).
+In all of this analysis we allow for heterogeneous treatment effects while still not relying on asymptotics (in contrast to some prior work on design-based instrumental variable analysis, such as Rosenbaum 1996).
+[^9] See Imbens and Rosenbaum (2005), Baiocchi et al.
+(2010), Keele et al.
+(2017), Kang et al.
+(2018), and Rambachan and Roth (2025) for additional prior work on design-based instrumental variable analysis.
 6.1 Setup
@@
 Suppose B1 (no defiers), B2 (relevance), and B3 (exact balance) hold. Then $\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}=\overline{\beta}_{1}(c)$.
-Proposition 2 shows that $\overline{\beta}_{1}(c)$ is point identified in finite populations under exact balance.
-In particular, it equals the finite population Wald estimand.
-This result is a finite population version of the Imbens and Angrist (1994) result, with one slight difference: The point identified parameter $\overline{\beta}_{1}(c)\coloneqq\left(\sum_{i=1}^{N}\beta_{i}\cdot\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)$ is a realized local average effect of treatment on the treated (LATT)—it is the average unit level causal effect among treated compliers (recall that $Z_{i}=X_{i}$ for compliers).
-In contrast, the population LATE is $\overline{\beta}(c)\coloneqq\left(\sum_{i=1}^{N}\beta_{i}\mathbbm{1}(T_{i}=c)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)$.
-The LATT is an ex ante random parameter, since the set of units who will be treated, and thus who appear in the parameter’s definition, depends on the realization of treatment assignment.
-This is analogous to Rosenbaum’s (2001) finite population analysis of ATT, where he noted that the ATT parameter is also ex ante random (he called the ATT the “attributable effect”).
-Another slight difference from the standard super-population analysis is that when there is one-sided noncompliance (so there are no always takers), the ATT and LATT are the same, but they do not equal LATE because there can be non-treated compliers.
-Finally, note that we can decompose LATE as $\overline{\beta}(c)=p_{1}(c)\overline{\beta}_{1}(c)+p_{0}(c)\overline{\beta}_{0}(c)$ where $p_{1}(c)\coloneqq\left(\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)\mathbbm{1}(Z_{i}=1)\right)/\sum_{i=1}^{N}\mathbbm{1}(T_{i}=c)$ is the proportion of units who are treated, among all compliers, and likewise for $p_{0}(c)$.
-So if we further assume that the average treatment effect for compliers in the instrument-on group is the same as for compliers in the instrument-off group—$\overline{\beta}_{1}(c)=\overline{\beta}_{0}(c)$—then the finite population Wald estimand equals LATE.
-This additional condition is analogous to part (iii) of B3.
+[…]
 6.3 Design-Based Sensitivity Analysis
@@
 $\left[\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}-\frac{K}{\pi},\;\frac{\overline{Y}_{z=1}-\overline{Y}_{z=0}}{\overline{X}_{z=1}-\overline{X}_{z=0}}+\frac{K}{\pi}\right].$
-The identified set in theorem 5 is analogous to the classical large population identified sets where instrument exogeneity is relaxed at the population level.
+The identified set in theorem 5 is analogous to the classical infinite population identified sets where instrument exogeneity is relaxed at the population level.
 For example, see Bound et al.
@@
 We conjecture that our analysis extends to this case but leave a full exploration to future work.
-7 Extensions
+7 Further Extensions
 7.1 Sampling
-From section 2.2 onward we assumed for simplicity that there was no sampling—all units in the population are observed, and the only uncertainty is about the unknown potential outcomes.
-Here we briefly discuss two extensions: An analysis of sampling by itself, and an analysis the combines both sampling and random assignment (e.g., as in Abadie et al.
+For most of this paper we assumed for simplicity that there was no sampling—all units in the population are observed, and the only uncertainty is about the unknown potential outcomes.
+Here we briefly discuss two extensions: An analysis of sampling by itself, and an analysis that combines both sampling and random assignment (e.g., as in Abadie et al.
 2020).
@@
 The same analysis can be done for the non-treated potential outcome, and they can be combined to do a design-based sensitivity analysis for the population ATE.
-Our approach treats sampling and randomization symmetrically—both lead to missing data which can be thought of as a finite population identification problem, and therefore analyzed using the tools of partial identification.
-However, there is a distinction between the two, which helps reconcile the difference between our definition 1 and definitions like Koopmans’ (1949) which assume away sampling uncertainty: Let $\mathcal{D}$ be the set of all designs, which are mappings from $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ into joint probabilities over sampling and treatment assignments, $(\mathbf{S},\mathbf{X})$.
+The Distinction Between Being Identified Versus Identifiable
+Our definition 1 treats sampling and randomization symmetrically—both lead to missing data which can be thought of as a finite population identification problem, and therefore analyzed using the tools of partial identification.
+This is perhaps unusual, since uncertainty due to sampling is traditionally separated from other types of uncertainty, as in Koopmans’ (1949, page 132) original definition of identification.
+Nonetheless, as sketched above, definition 1 provides a foundation to quantify uncertainty from both missing units and missing potential outcomes.
+That said, here we briefly discuss a distinction between the two, which helps reconcile the difference between our definition 1 and traditional definitions of identification in infinite populations like Koopmans’ (1949) which assume away sampling uncertainty.
+Consider the set of all designs, which are mappings from $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ into joint probabilities over sampling and treatment assignments.
 Say a parameter is point identifiable with respect to a set of assumptions $\mathcal{P}$ if there exists a design such that, for any $(\mathbf{Y}(1),\mathbf{Y}(0),\mathbf{W})$ satisfying the assumptions $\mathcal{P}$, and for any realization $(\mathbf{S},\mathbf{X})$ from that design, the subsequent identified set $\Theta_{I}$ is a singleton.
-Then the population mean $\overline{Y}$ is point identifiable with respect to A1, since we can in principle sample all units with probability 1.
+Then the population mean $\overline{Y}\coloneqq\frac{1}{N}\sum_{i=1}^{N}Y_{i}(X_{i})$ is point identifiable with respect to A1, since we can in principle sample all units with probability 1.
 But in any fixed dataset with $n<N$ it will only be partially identified.
 In contrast, ATE is not point identifiable with respect to A1 since we can never treat everyone and treat no-one.
+Thus, even though our definition of the identified set treats the ex post uncertainty due to missing units and missing potential outcomes symmetrically, our definition still allows these two kinds of uncertainty to be ex ante different, since one can in principle sample all units whereas we cannot even in principle treat all units.
 7.2 Variations on the Bounded Outcomes Assumption
 Since averages are influenced by outlier units, obtaining nontrivial bounds on ATE usually requires some kind of assumption that restricts the magnitude of outlier values.
-Throughout this paper we used a simple uniform bound $[y_{\text{min}},y_{\text{max}}]$ on potential outcomes, assumption A1.
+Throughout this paper we used a simple uniform bound $y_{\text{min}},y_{\text{max}}]$ on potential outcomes, assumption A1.
 First, it is important to recognize that this is a substantive identifying assumption, not a regularity condition—if these bounds are large then the researcher is explicitly allowing for large outliers and hence potentially large magnitudes of imbalance.
@@
 We use this version in one of our empirical applications, where the units are groups and the outcome depends on group size.
-(ii) One could impose a bounded unit level treatment effect assumption: $|Y_{i}(1)-Y_{i}(0)|\leq M$ for all $i\in\mathcal{I}$, where $M$ is a known sensitivity parameter.
+(ii) One could impose a bounded unit-level treatment effect assumption: $|Y_{i}(1)-Y_{i}(0)|\leq M$ for all $i\in\mathcal{I}$, where $M$ is a known sensitivity parameter.
 This assumption implies unit-specific bounds on the unobserved potential outcomes: $Y_{i}(x)\in[Y_{i}-M,Y_{i}+M]$.
-(iii) Alternatively, one could assume the sum of unit level treatment effect magnitudes is bounded: $\sum_{i=1}^{N}|Y_{i}(1)-Y_{i}(0)|\leq M$ for a known $M$.
+(iii) Alternatively, one could assume the sum of unit-level treatment effect magnitudes is bounded: $\sum_{i=1}^{N}|Y_{i}(1)-Y_{i}(0)|\leq M$ for a known $M$.
 This would allow for some units to have very large treatment effects, so long as not too many do.
@@
 This assumption implies that $Y_{i}(x)$ is in the interval $\frac{1}{N_{x}}\sum_{i=1}^{N}Y_{i}\mathbbm{1}(X_{i}=x)\pm 2\sqrt{N_{x}M}$.
-Hence it does not require users to specify a priori bounds $[y_{\text{min}},y_{\text{max}}]$; they only must choose $M$.
+Hence it does not require users to specify a priori bounds $y_{\text{min}},y_{\text{max}}]$; they only must choose $M$.
 (v) Finally, the covariate restrictions in section 5 can also be viewed as one way of restricting the magnitude of outliers.
@@
 It would be interesting to study how to extend our results to use the time series data as well, but we leave this to future work.
-The ATE point estimate is $13.7$ late parents, suggesting that the addition of a fee increased the number of late parents per week by about 14.
+The ATE point estimate is $13$ late parents, suggesting that the addition of a fee increased the number of late parents per week by 13.
 To quantify the uncertainty around this estimate, we conduct a design-based sensitivity analysis.
@@
 Moreover, even for large probabilities, the sets $\Theta_{I}(K(\alpha))$ are still mostly in the positive region.
-For example, the 90% set is about $[-6,22]$.
+For example, the 90% set is about [6,22][-6,22].
 There is at least a 90% chance that the true ATE is in this set.
@@
 Specifically, the outcome variable is the proportion of 38 management practices adopted in 2017, which was about 8 to 9 years after they received treatment.
-Since this outcome is a proportion, we set $[y_{\text{min}},y_{\text{max}}]=[0,1]$.
+Since this outcome is a proportion, we set $y_{\text{min}},y_{\text{max}}]=[0,1]$.
 (a) Top: $\Theta_{I}(K)$, bounds on ATE as a function of the maximum magnitude of ex post imbalance, $K$. Bottom: $\underline{p}(K)$.
@@
 Some of the 17 firms in the population have multiple plants.
-Treatment was assigned and administered at the plant-level, with at most 1 plant per firm treated.
+Treatment was administered at and varies at the plant-level.
 This implies that there are non-treated plants at firms with a different treated plant.
@@
 For simplicity we ignore all data from non-treated plants at treated-firms.
-This gives us a dataset where each unit is a single plant.
+We use the authors’ “experimental” dataset (as described in figure 1 of Bloom et al.
+2020), where each unit is a single plant.
 There are 11 treated plants and 6 control plants.
+[^10] See pages 16–17 of Bloom et al.
+(2013) for more details on the experimental design.
 The ATE point estimate is 0.13, suggesting that providing 4 months of support for changing management practices leads to a 13 percentage point increase in the proportion of practices that are still in place about one decade later.
@@
 2013)
-That is, to deal with the small population size they performed exact tests of the sharp null of no unit level treatment effects, assuming uniform randomization.
+That is, to deal with the small population size they performed exact tests of the sharp null of no unit-level treatment effects, assuming uniform randomization.
 In the 2013 paper they also used time series methods that are valid for fixed $N$ and large $T$, which we ignore since they cannot be used with the long run outcomes data.
@@
 Here we see that the sets all contain negative numbers for probabilities larger than 20%.
-For example, the 90% interval is $[-0.28,0.49]$.
+For example, the 90% interval is [0.28,0.49][-0.28,0.49].
 So there is at least a 90% chance that ATE is inside this interval, because this interval can be interpreted as a Bayesian credible set.
@@
 In this paper we studied identification in finite populations.
-We formally showed that, for any finite population size, randomization has no identifying power, because it does not guarantee any level of balance in unobservables.
-Nonetheless, we showed how to use randomization to derive empirical objective beliefs about the ex post level of balance.
-By combining finite population identified sets with the empirical objective worst case prior, we showed how to conduct a design-based sensitivity analysis, a new approach to quantifying uncertainty in finite populations.
-This approach yields uncertainty intervals that have (1) identification, (2) Bayesian, and (3) frequentist interpretations.
-From the frequentist perspective, our confidence intervals allow for arbitrary heterogeneous treatment effects, but do not rely on asymptotics.
-And we showed that our confidence intervals are consistent; they have statistical power.
-The key idea is that we are not doing inference based on hypothesis test inversion.
-Instead, we combine partial identification analysis with ideas from design-based inference to construct uncertainty intervals.
-Our approach also gave a new motivation for examining covariate balance, which can be used to shrink identified sets.
-We applied these ideas in both (1) classical randomized experiments and (2) the instrumental variable setting, focusing on one-sided noncompliance.
-Many open questions remain: In section 7 we sketched extensions to sampling, distributional measures of balance, and parameters beyond ATE, but left a full analysis for future work.
-We have focused on uniform randomization throughout this paper, but our approach likely generalizes to other types of randomization designs.
-One can also ask what optimal experimental designs are, given that the analyst will perform a design-based sensitivity analysis.
-Finally, we have focused on relatively simple settings with randomization by design, but an important next step is to extend this analysis to other settings like difference-in-differences or synthetic controls.
+We showed that the prior conventional definition of identification implies that every potential outcome for every unit is point identified under an overlap condition alone, even in observational settings where treatment depends on potential outcomes.
+Hence we argued that is not adequate for understanding what can be learned about treatment effects in real datasets.
+As our first main contribution, we developed a definition of identification that is not subject to this problem.
+Moreover, unlike the traditional design-based approaches to inference in finite populations, our definition of identification does not pre-suppose a known distribution from which treatment, an instrument, or a “shock,” was drawn.
+This allows it to be used in a wide variety of settings, including both experimental and observational studies.
+Building on our definition of identification, in our second main contribution we developed a new approach to quantifying uncertainty in finite populations, called a design-based sensitivity analysis.
+This approach combines design distributions and our finite population identified sets to construct new uncertainty intervals that have (1) identification, (2) robust Bayesian, and (3) uniform frequentist interpretations.
 Appendix A Proofs of Main Results
@@
 ∎
-Recall that $K^{\text{true}}(x)$ (eq (2)) denotes the true magnitude of imbalance in the $x$-potential outcome.
-Let $K^{\text{true,new}}(x)$ denote the same expression but with the random variables $X_{i}^{\text{new}}$ replacing $X_{i}$ for all $i\in\mathcal{I}$.
+Proof of proposition 1.
+Let $\mathbb{P}_{\mathcal{D}}$ denote the probability measure of $\mathcal{D}$.
+For any given $\mathbb{P}_{\mathcal{D}}$, there is a unique solution to $\mathbb{E}_{\text{design}}[m(\mathcal{D},\theta)]=0$, by assumption.
+Let $f(\mathbb{P}_{\mathcal{D}})$ denote this unique solution.
+This unique solution is assumed to equal $\theta^{\text{true}}$; hence $f(\mathbb{P}_{\mathcal{D}})=\theta^{\text{true}}$.
+Since $m$ is known, $f$ is known.
+Thus, by definition 2, $\theta^{\text{true}}$ is point identified.
+∎
+For any $N$-vector $\mathbf{A}=(A_{1},\ldots,A_{N})$, let
+$\text{DIM}(\mathbf{A},\mathbf{x}^{\text{new}})\coloneqq\frac{1}{N_{1}}\sum_{i=1}^{N}A_{i}\mathbbm{1}(x_{i}^{\text{new}}=1)-\frac{1}{N_{0}}\sum_{i=1}^{N}A_{i}\mathbbm{1}(x_{i}^{\text{new}}=0)$
+denote the difference-in-means of the $A_{i}$ variables across the treatment and control groups, when $\mathbf{x}^{\text{new}}$ is the assigned treatment vector.
+Then equation (2) can be written as $K^{\text{true}}(x)=|\text{DIM}(\mathbf{Y}(x),\mathbf{X})|$.
+Let $K^{\text{true,new}}(x)\coloneqq|\text{DIM}(\mathbf{Y}(x),\mathbf{X}^{\text{new}})|$.
 Proof of proposition 3.
@@
 Limits: Convergence to 1 as $K\rightarrow\infty$: Let $K\geq y_{\text{max}}-y_{\text{min}}$.
-Then A1 implies that for all dgps $\mathbf{Y}(1),\mathbf{Y}(0)$, for all realizations $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $|\overline{Y}_{1}(x)-\overline{Y}_{0}(x)|\leq K$, where $\overline{Y}_{g}(x)\coloneqq\frac{1}{N_{g}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=g)$ for $g\in\{0,1\}$.
+Then A1 implies that for all matrices of $\mathbf{Y}(1),\mathbf{Y}(0)$, for all realizations $\mathbf{X}$ of $\mathbf{X}^{\text{new}}$, $|\overline{Y}_{1}(x)-\overline{Y}_{0}(x)|\leq K$, where $\overline{Y}_{g}(x)\coloneqq\frac{1}{N_{g}}\sum_{i=1}^{N}Y_{i}(x)\mathbbm{1}(X_{i}=g)$ for $g\in\{0,1\}$.
 Thus $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=1$ for all such $K$.
@@
 We’ll show something stronger: There is an $\varepsilon>0$ such that $\underline{p}(K)=0$ for $K\in[0,\varepsilon)$.
-Recall that $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=\mathbb{P}(K^{\text{true,new}}(1)\leq K,K^{\text{true,new}}(0)\leq K)$.
+Recall that $p(K,\mathbf{Y}(1),\mathbf{Y}(0))=\mathbb{P}_{\text{design}}(K^{\text{true,new}}(1)\leq K,K^{\text{true,new}}(0)\leq K)$.
 And $\underline{p}$ is the infimum of this term over all $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ satisfying A1.
-Thus it suffices to find (i) a single dgp $\mathbf{Y}(1)$ and $\mathbf{Y}(0)$ that is consistent with the data and (ii) a small enough value of $K$ such that, for that dgp, just one of the two conditions $K^{\text{true,new}}(x)\leq K$ cannot hold.
-Consider the non-treated potential outcomes.
+Thus it suffices to find (i) a single choice of $(\mathbf{Y}(1),\mathbf{Y}(0))$ that is consistent with the data and (ii) a small enough value of $K$ such that, for that choice, $K^{\text{true,new}}(x)\leq K$ holds with probability zero for at least one $x\in\{0,1\}$.
+Without loss of generality, consider only $\mathbf{Y}(1)$; if we can find an $\varepsilon>0$ and a $\mathbf{Y}(1)$ that is consistent with the data and such that $K$-approximate mean balance in $\mathbf{Y}(1)$ is impossible for $K\in[0,\varepsilon)$, then it does not matter how the missing values of $\mathbf{Y}(0)$ are filled in, we will still have $\underline{p}(K)=0$ for $K\in[0,\varepsilon)$.
+This follows from $\mathbb{P}_{\text{design}}(K^{\text{true,new}}(1)\leq K,K^{\text{true,new}}(0)\leq K)\leq\mathbb{P}_{\text{design}}(K^{\text{true,new}}(1)\leq K)$.
 Order the observations so the first $N_{1}$ units are treated.
-We observe $\mathbf{Y}(0)_{1:N_{1}}$ in the data.
-It suffices to find values $\mathbf{Y}(0)_{N_{1}+1:N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ and an $\varepsilon>0$ such that $|(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{1:N_{1}})/N_{1}-(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{N_{1}+1:N})/N_{0}|>\varepsilon$, where $\boldsymbol{\iota}$ are vectors of 1’s.
-This can be done by choosing all components of $\mathbf{Y}(0)_{N_{1}+1:N}$ equal to $y_{\text{min}}$ or all components equal to $y_{\text{max}}$; note that in each case the average of the unobserved $Y_{i}(0)$’s is also equal to either $y_{\text{min}}$ or $y_{\text{max}}$.
-And note that $c\coloneqq(\boldsymbol{\iota}^{\prime}\mathbf{Y}(0)_{1:N_{1}})/N_{1}$ is a constant that is fixed in the data, with $c\in[y_{\text{min}},y_{\text{max}}]$.
-If $c=y_{\text{min}}$ then pick the unobserved $Y_{i}(0)$’s all equal to $y_{\text{max}}$ while if $c=y_{\text{max}}$ we pick the unobserved $Y_{i}(0)$’s all equal to $y_{\text{min}}$.
-In either case we set $\varepsilon=y_{\text{max}}-y_{\text{min}}>0$.
-If $c$ is strictly inside $[y_{\text{min}},y_{\text{max}}]$ then either choice for the $Y_{i}(0)$’s works; suppose we set them all to $y_{\text{min}}$.
-Then set $\varepsilon=|c-y_{\text{min}}|>0$.
+We observe $\mathbf{Y}(1)_{1:N_{1}}=\mathbf{Y}_{1:N_{1}}$ in the data.
+Suppose there exist values $\mathbf{Y}(1)_{(N_{1}+1):N}^{*}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ such that $\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{x}^{\text{new}})\neq 0$ for all $\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})$, where $\mathbf{Y}(1)^{*}\coloneqq(\mathbf{Y}_{1:N_{1}},\mathbf{Y}(1)_{(N_{1}+1):N}^{*})$.
+Let
+$\varepsilon\coloneqq\min_{\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})}\;|\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{x}^{\text{new}})|.$
+Since $\operatorname*{supp}(\mathbf{X}^{\text{new}})$ is finite, $\varepsilon>0$.
+Then for any $K\in[0,\varepsilon)$, the event $\{\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}}):|\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{x}^{\text{new}})|\leq K\}$ is the empty set.
+Hence $\mathbb{P}_{\text{design}}(|\text{DIM}(\mathbf{Y}(1)^{*},\mathbf{X}^{\text{new}})|\leq K)=0$ for all $K\in[0,\varepsilon)$.
+Consequently, the infimum over all possible completions $\mathbf{Y}(1)_{(N_{1}+1):N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ will also equal zero.
+Now suppose by way of contradiction that such a completion $\mathbf{Y}(1)^{*}$ does not exist.
+Then for all completions $\mathbf{Y}(1)_{(N_{1}+1):N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}$ there is some $\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})$ such that $\text{DIM}(\mathbf{Y}(1),\mathbf{x}^{\text{new}})=0$.
+That is,
+$y_{\text{min}},y_{\text{max}}]^{N_{0}}\subseteq\bigcup_{\mathbf{x}^{\text{new}}\in\operatorname*{supp}(\mathbf{X}^{\text{new}})}\left\{\mathbf{Y}(1)_{(N_{1}+1):N}\in[y_{\text{min}},y_{\text{max}}]^{N_{0}}:\text{DIM}\big((\mathbf{Y}_{1:N_{1}},\mathbf{Y}(1)_{(N_{1}+1):N}),\mathbf{x}^{\text{new}}\big)=0\right\}.$ (8)
+The right hand side of (8) is a finite union of hyperplanes, which therefore has empty interior, while the left hand side has non-empty interior since $y_{\text{min}}<y_{\text{max}}$.
+This is a contradiction.
 ∎
@@
 This implies that
-$\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K^{*}(\alpha)\text{ for $x\in\{0,1\}$})\leq\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K(\alpha)\text{ for $x\in\{0,1\}$}).$
+[…]
 Finally, by definition, $p(K,\mathbf{Y}(1),\mathbf{Y}(0))\coloneqq\mathbb{P}_{\text{design}}(K^{\text{true,new}}(x)\leq K\text{ for $x\in\{0,1\}$})$. Thus
@@
 Proof of theorem 4.
-[…]
+Part 1.
+We want to show that $\underline{p}(K^{\text{bp}})\xrightarrow{p}1$.
+There are two steps.
+First we use lemma LABEL:lemma:convergenceOfIdentifiedSet to show that the finite population breakdown point converges to a limiting breakdown point.
+Then we combine this step with lemma LABEL:lemma:underlinepGoesTo1 to get the final result.
+Suppose $\mu(1)-\mu(0)>0$; the less than zero case can be handled symmetrically.
   1. 1.
-[…]
+    Define the limit breakdown point $K_{\infty}^{\text{bp}}\coloneqq\sup\{K\geq 0:0\notin\Theta_{I,\infty}(K)\}$.
+Since $\Theta_{I,\infty}(K)$ is an interval and $\mu(1)-\mu(0)>0$, $K_{\infty}^{\text{bp}}$ can alternatively be written as the unique smallest solution to $\text{LB}_{K}^{\infty}(1)-\text{UB}_{K}^{\infty}(0)=0$ (defined in lemma LABEL:lemma:convergenceOfIdentifiedSet).
+Similarly, $K^{\text{bp}}\coloneqq\sup\{K\geq 0:0\notin\Theta_{I}(K)\}$ can be written as the unique smallest solution to $\text{LB}_{K}(1)-\text{UB}_{K}(0)=0$.
+Both the finite $N$ and limiting bound functions are continuous in $K$.
+And lemma LABEL:lemma:convergenceOfIdentifiedSet showed that the finite $N$ bound functions converge in probability to the limiting bound functions uniformly in $K$.
+Consequently, $K^{\text{bp}}\xrightarrow{p}K_{\infty}^{\text{bp}}$.
   2. 2.
-[…]
+    Since $\mu(1)-\mu(0)\neq 0$, $K_{\infty}^{\text{bp}}>0$.
+This combined with $K^{\text{bp}}\xrightarrow{p}K_{\infty}^{\text{bp}}$ implies that there is an $\varepsilon>0$ such that $\mathbbm{1}(K^{\text{bp}}\geq\varepsilon)\xrightarrow{p}1$.
+So $\underline{p}(K^{\text{bp}})=\underline{p}(K^{\text{bp}})\mathbbm{1}(K^{\text{bp}}<\varepsilon)+\underline{p}(K^{\text{bp}})\mathbbm{1}(K^{\text{bp}}\geq\varepsilon)=O_{p}(1)o_{p}(1)+\underline{p}(K^{\text{bp}})\mathbbm{1}(K^{\text{bp}}\geq\varepsilon)$.
+Finally, $1\geq\underline{p}(K^{\text{bp}})\mathbbm{1}(K^{\text{bp}}\geq\varepsilon)\geq\underline{p}(\varepsilon)\mathbbm{1}(K^{\text{bp}}\geq\varepsilon)$ by monotonicity of $\underline{p}$, and since, as a cdf, it is bounded above by 1.
+The last line converges in probability to 1 since $\varepsilon>0$ and by lemma LABEL:lemma:underlinepGoesTo1.
+Thus $\underline{p}(K^{\text{bp}})\mathbbm{1}(K^{\text{bp}}\geq\varepsilon)\xrightarrow{p}1$.
 Part 2. Fix $\alpha\in(0,1)$. Since $\Theta_{I}(K)$ is an interval, $\sup_{\theta\in\Theta_{I}(K(\alpha))}|\theta-\text{ATE}_{N}|$ equals
-$\max\left\{\left|\big{(}\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big{)}-\text{ATE}_{N}\right|,\;\left|\big{(}\text{LB}_{K(\alpha)}(0)-\text{UB}_{K(\alpha)}(1)\big{)}-\text{ATE}_{N}\right|\right\}.$
-[…]
+$\max\left\{\left|\big(\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big)-\text{ATE}_{N}\right|,\;\left|\big(\text{UB}_{K(\alpha)}(1)-\text{LB}_{K(\alpha)}(0)\big)-\text{ATE}_{N}\right|\right\}.$
+We will consider the first term only; the proof for the second term is analogous.
+We have $|\big(\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big)-\text{ATE}_{N}|\leq|\big(\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big)-\big(\mu(1)-\mu(0)\big)|+|\text{ATE}_{N}-\big(\mu(1)-\mu(0)\big)|$.
+The second term goes to zero by assumption.
+So consider the first term: $|\big(\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big)-\big(\mu(1)-\mu(0)\big)|\leq|\big(\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big)\allowbreak-\big(\text{LB}_{K(\alpha)}^{\infty}(1)-\text{UB}_{K(\alpha)}^{\infty}(0)\big)|+|\big(\text{LB}_{K(\alpha)}^{\infty}(1)-\text{UB}_{K(\alpha)}^{\infty}(0)\big)-\big(\mu(1)-\mu(0)\big)|$.
+The first term here goes to zero in probability by lemma LABEL:lemma:convergenceOfIdentifiedSet.
+The second term equals $|\mu(1)\rho+\max\{y_{\text{min}},\mu(1)-K(\alpha)\}(1-\rho)-\mu(0)(1-\rho)-\min\{y_{\text{max}},\mu(0)+K(\alpha)\}\rho-\mu(1)+\mu(0)|$.
+Recall that $K(\alpha)\coloneqq\inf\{K\geq 0:\underline{p}(K)\geq 1-\alpha\}$.
+So lemma LABEL:lemma:underlinepGoesTo1 implies that $K(\alpha)\rightarrow 0$ as $N\rightarrow\infty$, since $\alpha$ is strictly between 0 and 1.
+This implies that this second term goes to zero as $N\rightarrow\infty$.
+Thus we have shown that $\left|\big(\text{LB}_{K(\alpha)}(1)-\text{UB}_{K(\alpha)}(0)\big)-\text{ATE}_{N}\right|\xrightarrow{p}0$ as $N\rightarrow\infty$.
+∎
 Proof of lemma 1.